### v1

# Symmetry, Invariant Manifolds and Flow Reversals in Active Nematic Turbulence

Abstract
We investigate how symmetry, exact coherent structures (ECSs), and their invariant manifolds organize spontaneous flow reversals in a 2D active nematic confined to a periodic channel. In minimal flow units commensurate with the intrinsic active vortex scale, we use equivariant bifurcation theory to trace the origin of dynamically relevant ECSs via a sequence of symmetry-constrained local and global bifurcations. At low activity level, we identify relative periodic orbits, created via a sequence of SNIPER, homoclinic and heteroclinic bifurcations, whose invariant manifolds provide robust heteroclinic pathways between left- and right-flowing nearly uniaxial states. These result in several symmetry-dictated reversal mechanisms in the preturbulent regime, with and without vortex-lattice intermediate states. In the active turbulent regime, this ECS skeleton persists and organizes chaotic attractors exhibiting persistent two-way reversals. By classifying ECSs through their symmetry signatures, we relate a small set of ECSs embedded in turbulence back to the preturbulent branches, and show that typical turbulent trajectories repeatedly shadow these ECSs and their unstable manifolds, resulting in near-heteroclinic transitions between opposite-flow states. Our results establish that channel confined active nematic turbulence is organized by a low-dimensional, symmetry-governed network of invariant solutions and their manifolds, and identify dynamical mechanisms that could be exploited to design, promote, or suppress flow reversals in active matter microfluidic devices.
I Introduction
Active fluids serve as a paradigmatic class of non-equilibrium systems displaying a rich phenomenology, including spontaneous coherent flows, dynamical vortex patterns, and chaotic hydrodynamics (‘active turbulence’) [1, 2, 3]. Of particular recent interest among active fluids are active nematics (ANs), which are suspensions of rod-like units that convert energy into mechanical work at the individual level [4]. Examples of ANs include suspensions of elongated bacteria, kinesin-microtubule mixtures, and synthetic rod-like materials activated by chemical, optical or acoustic stimuli. These systems have attracted significant attention because they closely mirror biological processes from subcellular to multicellular scales and offer promising avenues for engineering applications.
Several continuum descriptions of active nematics have been derived using a variety of microscopic, phenomenological and data-driven techniques [5, 6, 7, 8]. The dynamic behavior of the continuum models is usually studied via a mix of numerical simulations [9, 10], and linear stability analysis of the analytically derivable steady ‘base’ states [11, 12, 6]. These base steady states are usually the simplest equilibria, or in certain cases, traveling wave solutions. Some works extend the analysis to predict periodic states [13, 14], and/or apply weakly nonlinear analysis [15, 16] to predict first few bifurcations from the base states [17]. Other recent works have explored the role of nonlinearity and non-reciprocity in characterizing vortical and traveling flow patterns in general active systems [18, 19, 20]. There have also been recent theoretical and computational efforts to derive coarse-grained statistical descriptions [3, 21, 22, 23] of active turbulence, focusing on energy spectra and spatial organization of statistically steady states, rather than deterministic dynamics.
While these studies provide invaluable insight into certain deterministic and statistical aspects of motion of active fluids, our understanding of the complexities of active turbulence in confinement remains limited [24, 25]. In this context, a deterministic, dynamical systems based approach to studying active turbulence was initiated in our prior works [26, 27]. The basic premise is that the infinite-dimensional phase space of an active fluid is populated by Exact Coherent Structures (ECS), which are exact solutions of the physical dynamics with distinct and regular spatiotemporal structure; examples include equilibria, periodic orbits, and traveling waves. The ECSs are connected by dynamical pathways in phase space, built of intersections of invariant manifolds. While original ideas go back to [28, 29], recent computational and algorithmic advances have resulted in successful application of ECS-based framework to partially resolve the long-standing problem of transition to turbulence in passive, high-Reynolds number fluid flows [30]. Specifically, convincing numerical and experimental evidence that turbulent trajectories track certain ‘dominant’ ECSs has been obtained for such passive,inertial systems [31].
A priori, there are multiple reasons to expect that ECS framework can be equally useful in understanding active turbulence by revealing the underlying, low-dimensional skeleton of the dynamics. First, robust motifs such as defect creation/annihilation events and vortical flows recur across experiments and simulations. Second, the lack of energy cascade in active turbulence [3] implies a narrower band of dynamically relevant scales as compared to inertial turbulence. By identifying the ECSs and heteroclinic connections that scaffold the turbulent attractor, one can: (i) clarify which nonlinear mechanisms are essential for active turbulence, (ii) provide a reduced description in terms of transitions between a small number of organizing states, and (iii) potentially control or design active flows via targeted perturbations that steer the system between selected coherent structures.
In our prior work [26], we analyzed the dynamics of the 2D AN periodic channel flow at a single, preturbulent activity level, identifying several dominant ECSs and the associated heteroclinic network. In the follow-up work [27], we extended this effort by computing a larger collection of ECSs spanning both preturbulent and turbulent activity levels in the same system, and showed numerically that typical trajectories in the preturbulent regime repeatedly shadow various ECSs before ultimately decaying into an attractor. However, the sheer number of ECSs present in the turbulent regime made it impossible to unambiguously detect shadowing by turbulent trajectories, leaving open the question of how ECSs and their invariant manifolds structure active turbulence. Moreover, the question of how the ECS repertoire is organized and how the different ECSs are related to one another was left unresolved.
Our present work is based on the key observation that, for a given geometry and boundary conditions, many continuum models of active nematics share the same symmetry group and the same action of this group on the relevant fields. Equivariant bifurcation theory then implies that the types of primary and secondary bifurcations, and the symmetries of their branches are dictated largely by the system symmetry and isotropy structure rather than by microscopic details of the model [32, 33]. Thus, a symmetry-based classification of ECSs and their bifurcations in one representative active nematic model can provide a template that is expected to be broadly applicable across different continuum descriptions (e.g. for microtubule–kinesin mixtures, bacterial suspensions, or actomyosin layers) that share the same symmetry class.
In the present work, we use equivariant bifurcation theory to systematically trace the origins of dynamically relevant ECSs such as dancing disclinations, oscillatory nearly unidirectional states, and sustained vortices, and to provide a symmetry-based organization of this ECS zoo within a minimal flow unit (MFU). The MFU is chosen to be large enough to support sustained active turbulence, yet small enough to permit a systematic ECS-based analysis. Guided by this framework, we investigate a robust and striking phenomenon reported in active nematic channel flow (and more generally active fluid) simulations and experiments: spontaneous flow reversals between nearly unidirectional flow states in active turbulence. We identify a phase space skeleton, consisting of ECSs and their heteroclinic connections, that captures the essential mechanisms behind flow reversals in the preturbulent regime. The relative simplicity of the MFU then allows us to extend this picture into the turbulent regime, where we uncover the corresponding organizing skeleton consisting of ECSs whose origin can be traced back to the preturbulent regime using their symmetry signatures. Finally, we provide direct numerical evidence that typical turbulent trajectories shadow these ECSs and their invariant manifolds in active turbulence.
In summary, our results show that active nematic turbulence in 2D channels is organized by a low-dimensional, symmetry-governed network of invariant solutions and manifolds, and we outline a general strategy for identifying similar symmetry-based mechanisms in other active turbulent flows.
The remainder of the paper is organized as follows. In Section II, we introduce the system model, summarize our previous ECS-based analysis of active nematic channel flow, and review experimental and theoretical work on flow reversals in active and passive fluids. In Section III, we use the symmetries of the 2D AN equations to carry out an equivariant bifurcation analysis in minimal flow units, tracing primary and secondary local and global bifurcations, and constructing a symmetry-organized catalog of dynamically relevant ECSs. In Section IV, we use this catalog to describe transient flow reversals in the preturbulent regime, identifying a small number of ECS families and their heteroclinic connections as the main mechanisms by which trajectories switch between opposite nearly unidirectional states. In Section V, we extend this picture into the active turbulent regime in two MFUs, characterize the chaotic and quasiperiodic sets that arise at high activity, and demonstrate that persistent two-way reversals are organized by a reduced ECS skeleton whose members can be related back to the preturbulent branches via their symmetry signatures.
II Model Description and Prior Work
II.1 Governing Equations
We consider the 2D Beris–Edwards continuum model for AN confined in a 2D periodic channel, a setup that has been used in several previous studies. [9, 34, 35, 26, 27]. It models the evolution of the velocity field $\mathbf{u}(\mathbf{x},t)=(u,v)$ and the nematic alignment tensor $\mathbf{Q}(\mathbf{x},t)$. The field $\mathbf{Q}$ is a traceless tensor defined as $Q_{ij}=q(n_{i}n_{j}-0.5\,\delta_{ij})$, where $q$ is the scalar order parameter describing the degree of nematic alignment, and the unit vector $\mathbf{n}$ denotes its direction.
The domain is taken to be a two-dimensional channel of size $0,\tilde{L}]\times[0,h]$, with $x\in[0,\tilde{L}$] the streamwise (horizontal) direction and $y\in[0,h]$ the wall-normal (vertical) direction. We impose no-slip boundary conditions on the velocity field $\mathbf{u}$ and perpendicular (homeotropic) anchoring on the nematic director $\mathbf{n}$ at the confining walls. We employ the nondimenzionalized equations obtained in [27]:
$\mathrm{Re}_{\mathrm{n}}\left(\partial_{t}+\mathbf{u}\cdot\boldsymbol{\nabla}\right)\mathbf{u}=-\boldsymbol{\nabla}p+2(\boldsymbol{\nabla}\cdot\mathbf{E})-\frac{\mathrm{R}_{\mathrm{a}}}{\mathrm{Er}}\left(\boldsymbol{\nabla}\cdot\mathbf{Q}\right),$
$\left(\partial_{t}+\mathbf{u}\cdot\boldsymbol{\nabla}\right)\mathbf{Q}+\mathbf{W}\cdot\mathbf{Q}-\mathbf{Q}\cdot\mathbf{W}=\lambda\mathbf{E}+\mathbf{H},$ (1)
$\boldsymbol{\nabla}\cdot\mathbf{u}=0,$
$\mathbf{H}=\mathbf{Q}-b\,\mathrm{Tr}\left(\mathbf{Q}^{2}\right)\mathbf{Q}+\nabla^{2}\mathbf{Q}.$ (2)
The boundary conditions are:
  • Velocity: No-slip at $y=0,h$, and periodic in $x$:
    $\mathbf{u}(x,0,t)=\mathbf{u}(x,h,t)=\mathbf{0},\quad\mathbf{u}(x+\tilde{L},y,t)=\mathbf{u}(x,y,t).$
  • Nematic tensor: Homeotropic anchoring with $q=1$ at $y=0,h$ and periodic in $x$:
    $\mathbf{Q}(x,0,t)=\mathbf{Q}(x,h,t)=\frac{1}{2}\begin{pmatrix}-1&0\\ 0&1\end{pmatrix},\quad\mathbf{Q}(x+\tilde{L},y,t)=\mathbf{Q}(x,y,t).$
In Eq. (1), $\mathbf{W}\cdot\mathbf{Q}-\mathbf{Q}\cdot\mathbf{W}$ captures the effect of flow rotation on the nematic director, and $\lambda\mathbf{E}$ is the flow-alignment term. The $\mathbf{H}$ term is the gradient of a free energy functional that promotes the creation of nematic order. We neglect the passive elastic stress which is subdominant compared to active and viscous stress. Here, $\mathrm{Re}_{\mathrm{n}}$ is the Reynolds number, $\mathrm{R}_{\mathrm{a}}$ the dimensionless activity, and $\mathrm{Er}$ the Ericksen number. Following [27, 4], we fix $\mathrm{Re}_{\mathrm{n}}=0.0136$ and $\mathrm{Er}=1$. As in the previous work [27], we also ignore the flow alignment to focus on the most relevant parts of the system and set the channel size to $\tilde{L}=80$ and $h=20$.
II.2 Exact Coherent Structure Framework for AN Channel Flow
In our previous works [26, 27], we recast the equations (1) as a dynamical system
$\frac{dX}{dt}=F(X;R_{a}),$ (3)
where $X=[\mathbf{u},\mathbf{Q}]$, the activity number $R_{a}$ is taken as the bifurcation parameter, and the associated flow map is given by $f^{t}(X_{0})=X_{0}+\int_{0}^{t}F[X(\tau)]d\tau$. We developed an open-source computational toolbox titled ‘Exact Coherent Strucures in Active Matter’ (ECSAct) [36] using the open-source pseudospectral code Dedalus [37] to compute ECSs and heteroclinic connections. The problem of computing each ECS is posed as a fixed-point equation (FPE), and these high-dimensional FPEs are solved via modified Newton-Raphson algorithms using adaptive ‘hookstep’ method and the matrix-free GMRES method [38].
In the first work [26], this computational framework was used to identify 46 ECSs (up to symmetry), including the unidirectional flow (UNI), the bidirectional flow (LAN), 11 unstable POs, and 33 RPOs in the preturbulent regime, see Fig. 1. We deem a parameter value to be preturbulent if there exists a regular (i.e., non-chaotic) attractor, i.e., a stable ECS in the phase space. Several heteroclinic connections between various ECSs were also computed, allowing the phase space to be organized as a directed graph with ECSs as nodes. The follow-up work [27] confirmed the previously reported general sequence of attractors that are observed while increasing activity: zero flow, unidirectional flow, vortex lattice, and the transition to turbulence. In the preturbulent regime, we identified several ‘shadowing events’, which are defined as finite time intervals over which a trajectory stays close to an ECS, and the trajectory and the ECS have quantitatively similar time evolution. While more than 200 ECS were found in the turbulent regime, no evidence of shadowing was found in that regime.
II.3 Flow Reversals in Active Nematics and Other Active Fluids
The Ph.D. thesis [39] reported experimental results on AN systems in long, narrow channels whose open ends were connected to an external chaotic AN reservoir. Persistent, nearly unidirectional flows were observed, and the direction of these uniaxial flow states (UNI) was selected at random via spontaneous symmetry breaking of the zero-flow (ZF) state, as expected. More importantly, spontaneous and repeated reversals were also observed between the left- and right-flowing UNI states, and the reversal frequency depended on the channel dimensions. The author used the channel-averaged streamwise velocity $\langle u\rangle$ as an order parameter to quantify these reversals. In these experiments, a typical reversal proceeded via an intermediate state: the UNI state with a parabolic velocity profile ($\langle u\rangle\approx\pm u_{0}$) first collapsed into a vortex-dominated configuration with negligible net flow ($\langle u\rangle\approx 0$), and then evolved into a UNI state whose mean flow was opposite in direction to the original one.
These reversals were attributed to a bend-instability mechanism. Starting from a nematic that is predominantly aligned along the streamwise direction, a bend instability will develop in the transverse direction [12] and reorient the director accordingly. The resulting transversely aligned state is, in turn, unstable to bend perturbations along the streamwise direction, leading to repeated switching between the two streamwise-aligned configurations. To capture this behavior statistically, the author derived and used a single-degree-of-freedom Langevin dynamics model with a double-well potential, fitted to the experimental data. The well minima corresponded to the two UNI states, and noise-driven fluctuations induced transitions between them.
AN flow reversals have also been observed in simulations of an (infinitely long) AN layer sandwiched between viscoelastic walls [14]. For sufficiently soft confinement, the authors found that the zero flow state undergoes a Hopf bifurcation to an oscillating flow state, in which the channel-averaged velocity reverses sign periodically in time, whereas stiffer walls produce steady unidirectional flow.
In polar active fluids, recent experimental, computational and analytical work [40] has argued that flow reversals are a precursor to active turbulence. The authors found that a dense suspension of swimming Bacillus subtilis confined to a cylindrical wells exhbit a single vortex steady flow at small radii. This state, which is analogous to the UNI in channel flow, loses stability as the radius is increased beyond a threshold. In the resulting oscillatory state, the vortex periodically reverses its direction. Using simulations and weakly nonlinear analysis of a Toner–Tu–Swift–Hohenberg model, it was shown that the onset of reversals is controlled by a global infinite-period bifurcation mediated by a resonant nonlinear interaction of the three lowest modes of the flow.
II.4 Possible Generic Dynamical Mechanisms Underlying Flow Reversals
Beyond active fluids, flow reversals have been extensively observed and studied in various settings [41, 42, 43, 44]. A prominent example is that of turbulent Rayleigh-Benard convection (RBC) [45], in which large-scale convection (LSC) reversals between the clockwise and anticlockwise vortical states are robustly observed [46]. In addition to numerical simulations, such reversals have been studied using both determninistic and stochastic reduced-order dynamical models [47, 42].
Regardless of their physical origins, systems that exhibit flow reversals are expected to share certain dynamical features. Typically, a quiescent base state first loses stability through spontaneous symmetry breaking, most naturally via a supercritical pitchfork bifurcation, producing two mirror-symmetric steady states. Exactly periodic reversals then imply the existence of a stable periodic orbit that alternately approaches these symmetry-related states, which may arise through a heteroclinic bifurcation or an infinite-period (global) bifurcation involving the two steady states. In turbulent regimes, where reversals are irregular and no simple periodic attractor exists, several possibilities arise. The turbulent trajectory may shadow a robust heteroclinic cycle [48, 49, 50, 51] built from the two symmetry-related states, possibly together with an intermediate state, although such cycles typically require strong symmetry constraints.
A more likely scenario is that turbulent trajectories are shadowing two-way heteroclinic connections between a pair of mirror-image unstable periodic orbits, each of which passes close to one of the steady states. Similar heteroclinic structures, in which each of two unstable periodic orbits passes near a different steady state and the two orbits are linked by two-way heteroclinic connections, have been found in numerous other low-dimensional settings [52, 53]. This reversal scenario could also involve additional ‘intermediate’ states and their invariant manifolds. This mechanism was shown to be the dynamical skeleton responsible for LSC reversals in a 2D forced Navier-Stokes system recently [54]. In that work, the author computed the two mirror-image asymmetric periodic orbits, and a symmetric ‘intermediate’ orbit. In a reduced phase space projection, turbulent trajectories are shown to shadow one asymmetric periodic orbit, then depart along its unstable manifold towards the intermediate periodic orbit, and finally either fall back to the original side, or cross-over towards the other asymmetric periodic orbit.
The main goal of the current work is to identify such dynamical structures organizing the reversals in the active turbulent regime, and establish their shadowing by typical turbulent trajectories. As a first step, the next section presents a local and global bifurcation analysis in the low-activity, preturbulent regime, clarifying how symmetry dictates the emergence of the most common ECS families in active nematic channel flows. The system naturally favors an intrinsic active vortex scale, resulting in threefold or fourfold symmetry of vortex lattice states in our chosen domain [27]. Hence, we carry out the equivariant analysis in shorter MFUs [55, 56], specifically $L=\tilde{L}/3$ and $L=\tilde{L}/4$, with a particular emphasis on the former due to its relevance in reversal dynamics.
III Local and Global Bifurcations in 2D AN Channel Flow: Role of Symmetry
The systematic way to perform bifurcation analysis of systems with symmetry is by using equivariant bifurcation theory [32, 57]. We begin this section by recalling some group theoretic definitions, and listing the symmetries of the AN 2D channel flow system. Then we revisit the equivariant analysis of primary bifurcations of the zero flow (ZF) state [33]. This prior work explains the emergence of UNI and LAN-like solutions in AN channel flows. We significantly extend the analysis of [33] to secondary local and global bifurcations, providing a partial but insightful explanation of the origin of dynamically relevant solutions and a classification based on their symmetry properties. Along the way, we highlight the consequences of symmetry-breaking, symmetry-preserving and symmetry-restoring bifurcations. A summary of the ECSs examined in this section, together with their symmetries, is provided in Table 5 in the appendix. Videos of the spatiotemporal evolution of these ECSs are provided in the Supplemental Material.
III.1 Preliminary definitions
A dynamical system of the form given by Eq. (3) is said be equivariant under a group action $\sigma$ if the time evolution and the group operation commute, i.e. for every state $X$ and time $T$, the relationship $\sigma(f^{T}(X))=f^{T}(\sigma X)$ is satisfied [58]. For an equivariant system, if $X$ is a solution, then $\sigma X$ is also a solution. Now, let the system Eq. (3) be equivariant under a group $G$, which implies that the above two properties are valid for all $\sigma\in G$.
We need to introduce additional definitions and notations from group theory that we use in the equivariant analysis. First, we denote $GL(\mathbb{R}^{n})$ as the group of all real invertible $n\times n$ matrices. The subgroup $O(n)\subset GL(\mathbb{R}^{n})$ is the set of matrices $A$ satisfying the orthogonality condition $A^{T}A=1$. We are primarily interested in groups that are isomorphic to $O(n)$ or to any closed subgroup of $O(n)$, i.e., compact Lie groups. Strictly speaking, the linear action of the symmetry $\sigma$ on the state $X$ is given by a matrix $\rho(\sigma)$, where the function $\rho$ is called a representation of $G$:
Definition 1 (Group Representation)
Let $G$ be a group and $V$ a vector space. A representation of $G$ on $V$ is a homomorphism
$\rho:G\to\mathrm{GL}(V),$
where $\mathrm{GL}(V)$ is the group of invertible linear transformations on $V$. For $\sigma\in G$ and $X\in V$, we will often write $\sigma X:=\rho(\sigma)X$.
If it is clear in the context, we will use the same symbol to denote the group element and the corresponding matrix.
We will say that a vector $X$ has $\sigma$ symmetry if it is invariant under the action of a group element $\sigma$, i.e., $\sigma(X)=X$. A solution of an equivariant system can have all, some or none of the symmetries $G$. Usually, we are interested in the group elements that act fix a specific vector $X$, motivating the following definition.
Definition 2 (Isotropy Group)
Let $G$ act on a set $V$. For an $X\in V$, the isotropy group (or stabilizer) of $X$ is the subgroup
$\Sigma_{X}=\operatorname{Stab}_{G}(X)=\{\sigma\in G\mid\sigma X=X\}.$
On the other hand, it is also relevant to know which vectors are invariant under the action of a specific subgroup.
Definition 3 (Fixed Point Subspace of a Subgroup)
Let $\Sigma\subset G$ be a subgroup. The fixed point subspace of $\Sigma$, denoted $\operatorname{Fix}(\Sigma)$, is defined as
$\operatorname{Fix}(\Sigma)=\{X\in V\mid\sigma X=X\text{ for all }\sigma\in\Sigma\}.$
An important property of fixed point subspaces is that they are flow invariant, i.e., if $X(0)\in Fix(\Sigma)$, then $X(t)\in Fix(\Sigma)$, for all $t>0$. The final definition we introduce in this section is that of the group orbit. Other useful definitions will be introduced as needed in the subsequent sections.
Definition 4 (Group Orbit)
For a point $X\in V$, the orbit of $X$ under the action of $G$ is the set
$GX=\{\sigma X\mid\sigma\in G\}.$
That is, the orbit of $X$ is the set of all points in $V$ to which $X$ can be moved by the action of elements of $G$.
III.2 Symmetries in the active nematics equations
The governing equations (1) with the imposed boundary conditions are equivariant under the continuous symmetries $\tau_{x}(\ell)$ corresponding to $x$-translation, the discrete symmetry $\sigma_{x}$ corresponding to reflection about the midplane $x$-axis, and the discrete symmetry $\sigma_{y}$ corresponding to reflection about the $y$- axis. The corresponding actions on the physical fields are:
$\sigma_{x}[u,v,Q_{11},Q_{12}](x,y)=[u,-v,Q_{11},-Q_{12}](x,h-y),$
$\sigma_{y}[u,v,Q_{11},Q_{12}](x,y)=[-u,v,Q_{11},-Q_{12}](L-x,y),$
$\tau_{x}(\ell)[u,v,Q_{11},Q_{12}](x,y)=[u,v,Q_{11},Q_{12}](x+\ell,y).$ (4)
Due to the periodic boundary conditions in $x$, the set of translations $\tau_{x}(\ell)$ is isomorphic to the rotational symmetry group $SO(2)$. The reflection symmetry $\sigma_{x}$ generates a cyclic group of order 2, which we denote as $\mathbb{Z}_{2}(\sigma_{x})$. Likewise, the reflection symmetry $\sigma_{y}$ generates the cyclic group denoted by $\mathbb{Z}_{2}(\sigma_{y})$. The x-translations and y-reflections do not commute, so they form $O(2)=SO(2)\rtimes\mathbb{Z}_{2}$ (a semdirect product). The x-reflection $\mathbb{Z}_{2}(\sigma_{x})$ commutes with this $O(2)$. Finally, the full symmetry group of the system is
$G=\mathbb{Z}_{2}(\sigma_{x})\times O(2).$
Aside from the work of Marzorati and Turzi [33], which we discuss in the next subsection, prior stability and bifurcation studies of active nematic channel flows have not systematically exploited these symmetries. This omission obscures the underlying structure of the infinite-dimensional dynamical system [59]. In this work, we use equivariant bifurcation theory to systematically predict the emergence of new steady and time-dependent solution branches and to determine the symmetries of these solutions.
Some solutions of Eq. (1) have already been classified by their symmetries previously [26]. The ZF inherits all the symmetries of $G$, but UNI and LAN lack the $\sigma_{y}$ and $\sigma_{x}\sigma_{y}$ symmetries, respectively. More generally, the plethora of ECSs found in [26] exhibit a rich variety of patterns with observable spatiotemporal symmetries (see Fig. 1). The system selects preferable discrete translation symmetries to form vortex solutions [26, 27], many of them corresponding to oscillatory solutions which maybe traveling waves, periodic or relative periodic orbits. In the rest of this section, we use equivariant bifurcation theory to identify the origins of key ECSs, organize the associated solution branches by their shared symmetries, and assemble a partial isotropy lattice that connects symmetry-distinct branches through sequences of symmetry-breaking bifurcations.
Figure 1: The velocity field snapshots of some ECSs at representative activity numbers $\mathrm{R_{a}}=1.5$ and $\mathrm{R_{a}}=4.5$ are shown here with their symmetries in the full-length channel, $L=\tilde{L}$. The color map indicates vorticity.
III.3 Primary bifurcation in active nematics
The generalized equivariant branching lemma [57, 58, 32] described below is a fundamental tool in equivariant bifurcation analysis, and provides the conditions for the existence of a symmetric fixed point branch emerging from a singular fixed point $X^{*}$ at a bifurcation value $R_{a}=R_{a}^{*}$. To apply this framework to our high-dimensional system (Eqs. 3), we restrict the system to the center manifold, where the evolution is given by $\dfrac{dY}{dt}=F_{c}(Y,\mu)$, where $\mu=R_{a}-R_{a}^{*}$, and the system satisfies $F_{c}(0,0)=0$, and $D_{Y}F_{c}(0,0)=0$. The reduced system $F_{c}$ is $\Gamma$-equivariant, where $\Gamma$ is the isotropy group of $X^{*}$, i.e., $\Gamma=\{\sigma\in G|\sigma(X^{*})=X^{*}\}$.
Theorem 1 (Generalized Equivariant Branching Lemma)
Let $F_{c}$ be as defined above, and let $\Sigma$ be a isotropy subgroup of $\Gamma$. Assume that $D_{\mu}F_{c}(0,0)w\neq 0$ for nonzero $w\in\operatorname{Fix}(\Sigma)$, and $Fix(\Gamma)=0$. Then, if $\dim(\operatorname{Fix}(\Sigma))=1$ (axial subgroup), there exists a smooth branch of fixed points of $F_{c}$ (and hence, of $F$) emerging from the origin with symmetry $\Sigma$.
Note that if $X^{*}$ has all the symmetries of $G$, then $\Gamma=G$, but otherwise, $\Gamma$ is a proper subgroup of $G$. A key observation is that $\Gamma$ commutes with the Jacobian $L\coloneq D_{X}F(X^{*},R_{a}^{*})$, which implies that $\Gamma(\ker L)=\ker L$, i.e., $\Gamma$ maps the kernel (or the center subspace) back to itself. This fact allows us to study the bifurcation problem via the reduced system $F_{c}$ by analyzing the action of $\Gamma$ on the center subspace.
Using the EBL, Marzorati and Turzi [33] performed an equivariant bifurcation analysis of the 2D AN channel system to obtain solutions bifurcating from the zero-flow (ZF) state. We discuss their calculations of this primary bifurcation to illustrate the application of the EBL and set the stage for calculation of secondary bifurcations that follow.
By restricting the analysis to (x-) translation invariant solutions, they only considered equivariance with respect to the Klein four-group $\hat{G}=\mathbb{Z}_{2}(\sigma_{x})\times\mathbb{Z}_{2}(\sigma_{y})$. The ZF solution has the full symmetry, and hence $\Gamma=\hat{G}$. The kernel of the Jacobian at ZF has two dimensions that we associate to $\mathbb{R}^{2}$ spanned by the two eigenvectors corresponding to zero eigenvalues. In this subspace, the matrix representations of the group $\Gamma=\{I,\sigma_{x},\sigma_{y},\sigma_{x}\sigma_{y}\}$ were computed in the eigenvector basis, and found to be
$I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\quad\sigma_{x}=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},\quad\sigma_{y}=\begin{bmatrix}-1&0\\ 0&-1\end{bmatrix},\quad\sigma_{y}\sigma_{x}=\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}$.
The four proper subgroups of $\Gamma$ are $\{I\}$, $\mathbb{Z}_{2}(\sigma_{x})$, $\mathbb{Z}_{2}(\sigma_{y})$, and $\mathbb{Z}_{2}(\sigma_{x}\sigma_{y})$. The corresponding fixed point subspaces are
$\text{Fix}(\{I\})$ $=\mathbb{R}^{2}\text{ : 2D fixed point subspace,}$
$\text{Fix}(\mathbb{Z}_{2}(\sigma_{x}))$ $=\{(a,0)\in\mathbb{R}^{2}\mid a\in\mathbb{R}\}\text{ : 1D fixed point subspace},$
$\text{Fix}(\mathbb{Z}_{2}(\sigma_{y}))$ $=\{(0,0)\}\text{ : 0D (trivial) fixed point subspace},$
$\text{Fix}(\mathbb{Z}_{2}(\sigma_{x}\sigma_{y}))$ $=\{(0,b)\in\mathbb{R}^{2}\mid b\in\mathbb{R}\}\text{ : 1D fixed point subspace.}$
Hence, there are two axial (i.e. with 1D fixed point subspaces) subgroups $\Sigma$, and by the EBL, we have two pitchfork bifurcations. The first bifurcation leads to UNI-like solutions with reflection symmetry $\Sigma_{1}=\sigma_{x}$, and the second to LAN-like solutions with symmetry $\Sigma_{2}=\sigma_{x}\sigma_{y}$. Due to the pitchfork nature of these bifurcations, each new solution branch has a symmetric counterpart obtained by applying $\sigma_{y}$. Therefore, a total of four nontrivial solution branches emerge.
Building upon this result, we carry out secondary bifurcation analysis from the UNI and LAN states in the following sections.
III.4 Secondary bifurcations from UNI
The UNI solution loses stability when a (‘first’) pair of complex eigenvalues crosses the imaginary axis. The corresponding eigenvectors break translation symmetry, and have isotropy group $\Sigma_{p_{1(\text{U})}}=\operatorname{gen}\{\sigma_{x}\tau_{x}(L/2)\}$, corresponding to the ‘shift-reflect’ symmetry [59]. Here, we use the notation $p_{1(\text{U})}$ to denote the first complex pair of eigenvalues (and the corresponding eigenspace) of the UNI. The imaginary nature of the eigenvalues suggest the possibility of a Hopf bifurcation. This bifurcation is expected to lead to oscillatory solutions that respect the underlying symmetry $\Sigma_{p_{1(\text{U})}}$. Such oscillatory states are typically small perturbations of the UNI configuration, and have been reported in previous studies [34, 60, 61, 9, 62]. Here, we explore a scenario in which traveling waves emerge from this destabilization of UNI.
Secondary bifurcations are often more challenging to analyze than primary bifurcations, as the standard equivariant bifurcation theorems typically do not apply directly. This difficulty arises when a solution undergoing bifurcation has less symmetry than the full group $G$, and a resulting continuous group orbit. The presence of the latter can dramatically modify the bifurcation scenario. While UNI (and LAN) have a smaller symmetry than $G$, their group orbits are discrete, and consist of two points each (the two y-reflections via $\sigma_{y}$). This allows the direct use of standard equivariant bifurcation theory.
III.4.1 Traveling waves emerge from UNI
We now state a version of the equivariant Hopf theorem [58, 32, 57] (EHT) which will be used to predict bifurcating oscillatory solutions with symmetry at a singular fixed point with a pair of purely pair of eigenvalues. As in the EBL case, we consider the reduced-dynamics on this 2D center subspace, which we identify as $\mathbb{R}^{2}$ spanned by the two eigenvectors. The reduced system $\dot{Y}=F_{c}(Y,\mu)$ is $\Gamma$-equivariant, where $\Gamma$ is the isotropy subgroup of the fixed point $X^{*}$. Let the eigenvalues be $\lambda(\mu)=\sigma(\mu)\pm i\rho(\mu)$, with bifurcation occurring at $\mu=0$.
Theorem 2
Let $F_{c}$ be as defined above, and $\Gamma$ act simply on the center subspace. Assume $\frac{d\sigma}{d\mu}|_{(\mu=0)}$ is different from zero. Suppose there exists an isotropy subgroup $\Sigma\subset\Gamma\times SO(2)$ such that $dim(Fix(\Sigma))=2$. Then there exists a unique branch of time-periodic solutions with period $\approx 2\pi/\omega$ emerging from the origin with spatio-temporal symmetry $\Sigma$, where $\omega=\rho(0)$.
To apply this to bifurcation of UNI, we consider the isotropy subgroup of UNI as the base group, denoted by $\Gamma=\mathbb{Z}_{2}(\sigma_{x})\times\mathrm{SO}(2)\subset G$. According to EHT, we seek a subgroup $\Sigma\subset\Gamma\times\mathrm{SO}(2)$ that fixes pointwise a two-dimensional subspace, which in this case is the entire $\mathbb{R}^{2}$.
Next, we obtain the representation of $\Gamma=\mathbb{Z}_{2}(\sigma_{x})\times\mathrm{SO}(2)$ on $\mathbb{R}^{2}$ in the eigenvector basis. On this center subspace, we find that refection $\sigma_{x}$ acts as
$\sigma_{x}=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}$, (5)
and the translations act as the rotation matrices
$R(\phi)=\begin{pmatrix}\cos\phi&-\sin\phi\\ \sin\phi&\cos\phi\end{pmatrix},\quad\phi=\frac{2\pi\ell}{L},\quad\ell\in[0,L),\quad\phi\in[0,2\pi),$ (6)
where the angle $\phi$ parametrizes the spatial rotation. The representation of the temporal group is
$R_{t}(\theta)=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix},\quad\theta\in[0,2\pi),$ (7)
where $\theta$ parametrizes the ‘rotation’ in time.
The representation of $\Gamma$ is irreducible but not absolutely irreducible on $\mathbb{R}^{2}$, i.e., it acts simply on $\mathbb{R}^{2}$. Indeed, the only subspaces invariant under the action of all group elements given by Eqs. (5) and (6) are the origin, and the whole space $\mathbb{R}^{2}$, and these matrices commute with each other.
The spatiotemporal isotropy subgroup $\Sigma$ that fixes $\mathbb{R}^{2}$ (pointwise) consists of all group elements whose representation is the identity matrix, which yields:
$\Sigma=\{I,\sigma_{x}R(\pi),\sigma_{x}R_{t}(\pi),R(\theta)R_{t}(\theta)\}.$ (8)
Recall that $R(\pi)$ is the $L/2$ spatial shift, and hence the purely spatial discrete symmetry $\sigma_{x}R(\pi)$ is nothing but the shift-reflect symmetry $\Sigma_{p1(U)}$. As rotation in time $R_{t}(\pi)$ has the same representation as the spatial $L/2$ shift $R(\pi)$, the discrete spatiotemporal symmetry $\sigma_{x}R_{t}(\pi)$ is also the shift-reflect symmetry. Finally, the group elements of the form $R(\theta)R_{t}(\theta)$ describe continuous spatiotemporal symmetries in which each shift (or rotation) in space is compensated by a corresponding rotation in time.
Then, the EHT says that a unique branch of small-amplitude time-periodic solutions bifurcates from the UNI state, with spatiotemporal symmetry group $\Sigma$. Concretely, $R(\theta)R_{t}(\theta)$ symmetry implies the solutions are (small-amplitude) traveling (or equivalently, rotating) waves, while the other two symmetries require that these traveling waves possess the purely spatial shift-reflect symmetry $\sigma_{x}\tau_{x}(L/2)$.
Figure 2: Bifurcation diagram showing TW speed vs $\mathrm{R}_{\mathrm{a}}$ for the MFU with length $L=\tilde{L}/4$. As UNI loses stability at $\mathrm{R}_{\mathrm{a}}\approx 0.6735$, a stable TW emerges through a Hopf bifurcation. An unstable TW exists prior to this value and collides with the stable TW in a saddle-node bifurcation at $\mathrm{R}_{\mathrm{a}}\approx 0.67561$, leading to their mutual disappearance.
Figure 3: A snapshot of the nematic director and velocity fields of the stable traveling wave at $\mathrm{R}_{\mathrm{a}}=0.67558$ (for MFU with $L=\tilde{L}/4$). This solution lies in the fixed-point subspace of $\Sigma_{p_{1(\text{U})}}$, i.e., it has the shift-reflect symmetry.
Figure 4: Projection onto the reduced phase space ($\langle Q_{11}\rangle$, $\langle v^{2}\rangle$), showing the stable and unstable TWs before the SNIPER bifurcation, with a heteroclinic connection passing close to UNI. Note the RPO that emerges after SNIPER has a shape similar to that of the vanished heteroclinic network.
Figure 5: Inverse-scaling-law fit for the period along the emerging RPO branch., supporting the SNIPER bifurcation scenario. Here, $A\approx 35.21$ and $B\approx-20.683$. Notice that the fitted bifurcation value $\mathrm{R}_{\mathrm{{a_{c}}}}\approx 0.674$ is quite close to the value $\mathrm{R}_{\mathrm{{a_{c}}}}\approx 0.67558$ obtained from the calculations based on data before the bifurcation.
Figure 6: Saddle-node infinite-period bifurcation of traveling waves following the Hopf bifurcation of UNI equilibrium in the MFU with $L=\tilde{L}/4$.
III.4.2 Saddle-node bifurcation of traveling waves
The equivariant bifurcation analysis so far is general enough to be applied independently of the channel dimensions. For a MFU of length $L=\tilde{L}/4$, our computations yield the stable TW predicted in the previous section (see Fig. 6). This TW exists in the range $\mathrm{R}_{\mathrm{a}}\approx 0.6735$ to $\mathrm{R}_{\mathrm{a}}\approx 0.67561$, where it collides with an unstable traveling wave (UTW) and disappears in a saddle-node bifurcation. This is a bifurcation from the TW solution, a continuous group orbit. As remarked previously, we need to use an extension of the standard equivariant bifurcation theory that accounts for the effects of the continuous symmetries [63] of the TW solution, in order to theoretically confirm this bifurcation.
For a $\Gamma$-equivariant dynamical system, suppose the group orbit $\Gamma X_{0}$ of a point $X_{0}$ has dimension at least one (i.e., it is not a discrete set of points), and let $\Sigma_{X_{0}}$ be the isotropy subgroup of $X_{0}$. This group orbit can be flow-invariant (even if $X_{0}$ is not an equilibrium) when the dynamics of the system cause the solution to evolve purely along the group orbit, in which case the group orbit is referred to as a relative equilibrium. The TW and UTW solutions are both examples of such relative equilibria.
The bifurcation analysis of a relative equilibrium begins by locally decomposing the vector field $F$ into the sum of a normal (to the group orbit) component and a tangential (i.e., along the group orbit) component: $F=F_{N}+F_{T}$, near the relative equilibrium $\Gamma X_{0}$ [63]. Roughly speaking, the dynamics of $F_{N}$ can be understood as the dynamics of $F$ modulo the drift along the group orbit. To analyze the bifurcation, we consider the restriction $H(X,\mu)$ of $F_{N}(X,\mu)$ to the subspace $N_{X_{0}}$, which is normal to the group orbit at the point $X_{0}$. A classical equivariant bifurcation analysis can then be performed on $H$. The following proposition, which is an extension of EBL, is needed to address the case where $\Sigma_{x_{0}}$ acts trivially on a 1D or 2D center subspace [63].
Proposition 1
Suppose the action of the isotropy subgroup $\Sigma_{X_{0}}$ on the eigenspace is trivial.
  • If $D_{X}H(X_{0},0)$ has a zero eigenvalue, then generically $H$ undergoes a limit point (saddle-node) bifurcation.
  • If $D(X_{0},0)$ has a pair of purely imaginary eigenvalues $\pm i\omega$, then generically $\pm i\omega$ are simple and $H$ undergoes a Hopf bifurcation to a unique periodic solution.
For the case at hand, we approximate the relevant eigenspace by computing the subspace corresponding to the eigenvalue with smallest non-zero real part for linearization around the UTW, close to the saddle-node bifurcation point. As expected, this space is one-dimensional, associated with a small real eigenvalue, and is invariant under the symmetry $\sigma_{x}\tau_{x}(L/2)$. This implies that the action of the isotropy subgroup $\Sigma_{x_{0}}$ on the eigenspace is trivial, indicating that no symmetry is broken at the bifurcation. By Proposition 1, we conclude that the bifurcation is a saddle-node bifurcation of relative equilibria. This result is consistent with our numerical finding that the two branches of traveling waves coalesce at the bifurcation point.
The rigorous dimension count of the group orbits of equilibria and periodic orbits is given by the following theorem.
Theorem 3
[63] Consider a $\Gamma$- equivariant dynamical system.
  • Let $X$ be an equilibrium of the normal vector field, with isotropy group let $\Sigma_{X}$. Then the group orbit of $X$ is generically a quasiperiodic solution with $k$ basic frequencies, where $k=\operatorname{rank}\left(N(\Sigma_{X})/\Sigma_{X}\right).$
  • Let $Y$ be a periodic orbit of the normal vector field, and $\Sigma_{Y}$ be the group of symmetries mapping $Y$ to itself. Then the group orbit $\Gamma Y$ is generically a quasiperiodic solution with $k$ basic frequencies, where $k=\operatorname{rank}\left(N(\Sigma_{Y})/\Sigma_{Y}\right)+1$.
    Here, $N(\Sigma_{X})$ denotes the normalizer of $\Sigma_{X}$ in $\Gamma$, and rank refers to the maximum dimension of the resulting torus.
Remark: Under the bifurcation scenario of Proposition 1, the resulting equilibria ($X$) and periodic orbits ($Y$) bifurcating from $X_{0}$ both inherit the symmetry of $X_{0}$. Hence in such situations, the Theorem 3 dimension counts can be replaced by $k=\operatorname{rank}\left(N(\Sigma_{X_{0}})/\Sigma_{X_{0}}\right)$, and $k=\operatorname{rank}\left(N(\Sigma_{X_{0}})/\Sigma_{X_{0}}\right)+1$, respectively.
Using the dimension count formula of Theorem 3 with $\Gamma=\mathbb{Z}_{2}(\sigma_{x})\times\mathrm{SO}(2)$ (inherited from UNI), and isotropy group being the shift-reflect $\Sigma_{X_{0}}=\mathbb{Z}_{2}(\sigma_{x}\tau_{x}(L/2))$, we get $k=1$. This is consistent with our observation that the bifurcating solutions are traveling waves, where time evolution coincides with translations $\tau_{x}(l)X$ ($0\leq l\leq L$).
III.4.3 Saddle-node Infinite-period bifurcation and the origin of homoclinic-like relative periodic orbits
The saddle-node bifurcation of TWs discussed above is part of a global bifurcation scenario involving a saddle-node infinite-period bifurcation (SNIPER), as illustrated in Fig. 6. In the interval $0.6737\leq\mathrm{R}_{\mathrm{a}}\leq 0.67561$, UNI, TW, and UTW coexist, with evidence of heteroclinic connections between them. To investigate whether transversal or tangential intersections among their manifolds are possible, we focus on the activity value $\mathrm{R}_{\mathrm{a}}=0.675$. Generically in an NN- dimensional space, an unstable manifold of dimension $n_{u}$ will intersect a stable manifold of dimension $n_{s}$ in $d=n_{u}+n_{s}-N$ dimensional submanifold [64].
UNI and UTW each have two-dimensional unstable manifolds, which allows for transversal intersections with the stable manifold of TW. However, a transversal connection between the unstable manifold of UTW and the stable manifold of UNI is not possible. Perturbing UNI along its unstable manifold leads TW, while perturbing UTW in an unstable direction also drives the system toward TW. Interestingly, we also found a heteroclinic connection from UTW to UNI very close to $\mathrm{R}_{\mathrm{a}}=0.675$. The various connections are illustrated in Fig. 6.
As the activity increases, the two traveling waves get closer and disappear in the saddle-node bifurcation at $\mathrm{R}_{\mathrm{a}}\approx 0.6756$. Beyond this point, a family of relative periodic orbit (RPO) with isotropy group $\Sigma_{p_{1}(U)}$ emerges, passing very close to UNI. The shape of the resulting RPOs, shown in the reduced phase space in Fig. 6, closely resembles the vanished heteroclinic network, and their periods follow an inverse law characteristic of a SNIPER bifurcation [65] (see Fig. 6).
In MFU with length $L=\tilde{L}/3$, the predicted traveling wave (TW) bifurcating from UNI proved elusive to detect. We only found an unstable TW that disappears when UNI loses stability and can be traced back to lower activity numbers, around $\mathrm{R}_{\mathrm{a}}\approx 0.55$. In this MFU, the branch of bifurcating RPOs originating near $\mathrm{R}_{\mathrm{a}}\approx 0.71$ also exhibits decreasing periods as activity increases, consistent with a SNIPER bifurcation law (see Fig. 7). We therefore conclude that the saddle-node bifurcation likely occurs near $\mathrm{R}_{\mathrm{a}}\approx 0.71$ in the MFU with $L=\tilde{L}/3$, although the small-amplitude stable TW is difficult to detect due to the close proximity (in parameter space) of the Hopf and saddle-node bifurcations.
Due to their proximity to UNI in phase space, we refer to these RPOs as homoclinic-like RPOs characterized by recurrent processes of defect creation and annihilation. We refer to the RPO family in Fig. 7 as HRPOT1a, where $Tk$ in the subscript implies the $L/k$ translation symmetry, and the following letter distinguishes different families $(a,b,\dots)$. Since the dynamics leading to flow reversals are closely associated with the exact coherent structures (ECSs) that emerge in the channel with $L=\tilde{L}/3$, we henceforth restrict our analysis to this domain (and subdomain $L=\tilde{L}/6$) for the remainder of the paper. The MFU with $L=\tilde{L}/3$ is henceforth referred to as MFU A.
(a) Reduced phase space projection of four members of the homoclinic-like RPO family HRPOT1a at $\mathrm{R_{a}}=0.72$, $0.90$, $1.02$, and $1.24$, shown in red, blue, black, and green solid curves, respectively. Their $\sigma_{y}$ counterparts are shown as dashed curves. For HRPOT1a at $\mathrm{R_{a}}=1.24$, we show velocity snapshots at two locations along the orbit: one corresponding to a vortex-dominated flow state and the other a nearly unidirectional flow state.
(b) Inverse law scaling fit between the period and the activity. Here, $A\approx 40.399$, $B\approx-16.209$, and $\mathrm{R_{a_{c}}}\approx 0.702$.
Figure 7: Confirmation that homoclinic-like RPOs (of the HRPO1a family) originate from a SNIPER bifurcation in the MFU A ($L=\tilde{L}/3$).
III.4.4 From Homoclinic-like RPOs to Heteroclinic-like POs
As activity increases beyond the SNIPER value, the periods of homoclinic-like RPOs initially decrease (as predicted by the inverse law), but eventually the RPOs start to become larger again, until the RPO branch collides with UNI and disappears through a homoclinic bifurcation at approximately $\mathrm{R}_{\mathrm{a}}\approx 1.38$. Our computations confirm the characteristic logarithmic scaling associated with this type of bifurcation (see Fig. 8b).
Immediately beyond this point, a branch of heteroclinic-like periodic orbits emerges, passing very close to both UNI and $\sigma_{y}\texttt{UNI}$ (see Fig. 8a), via a heteroclinic bifurcation involving these equilibria. For these activity numbers, the period data again fits well to a logarithmic curve, as shown in Fig. 8c. The solutions on this new branch of periodic orbits are denoted as belonging to the HTPOT1a family, where we use the same naming convention as before. The origin of these heteroclinic-like POs is a type of symmetry-increasing bifurcation [66, 67] in which the two RPOs, related by the $\sigma_{y}$ symmetry, disappear to create a PO whose orbit is $\sigma_{y}$- symmetric . These kinds of solutions are important because they contain trajectories connecting the two sides of the reduced phase space, i.e., $\langle u\rangle>0$ and $\langle u\rangle<0$. The HTPOT1a members are additionally invariant under the symmetry $\sigma_{x}\tau_{x}(L/2)$ inherited from the HRPOs.
(a) Two homoclinic-like RPO members of the HRPOT1a family at $\mathrm{R}_{\mathrm{a}}=1.35\text{ and }1.37$ shown as solid red and blue curves, respectively, with their $\sigma_{y}$ symmetry-related counterparts shown as dashed curves. And two heteroclinic-like PO members of the HTPOT1a family at $\mathrm{R}_{\mathrm{a}}=1.39\text{ and }1.41$ shown as black and green curves, respectively. These POs are also pre-periodic orbits with respect to $\sigma_{y}$, so their orbits are invariant under this symmetry. The black arrow shows the direction of increasing activity ($\mathrm{R}_{\mathrm{a}}).$
(b) Logarithmic scaling of period near homoclinic bifurcation in which the HRPOs disappear. Here, $A\approx 11.515$, $B\approx 40.69$, and $\mathrm{R}_{\mathrm{a_{c}}}\approx 1.381$.
(c) Logarithmic scaling of period near heteroclinic bifurcation after which the HTPOs emerge. Here, $A\approx 33.865$, $B\approx-0.755$, and $\mathrm{R}_{\mathrm{a_{c}}}\approx 1.38$.
Figure 8: Two consecutive global bifurcations, both with logarithmic scaling consistent with homoclinic/heteroclinic bifurcations. The first (homoclinic) bifurcation results in the vanishing of homoclinic-like RPOs. The second (heteroclinic) bifurcation leads to emergence of heteroclinic-like POs. In both cases, the RPOs and POs collide with the UNI (and $\sigma_{y}$ UNI) solution at the respective bifurcation points. Note that the estimated bifurcation values $\mathrm{R}_{\mathrm{a_{c}}}$ are very close to each other.
We find another heteroclinic bifurcation again involving UNI and $\sigma_{y}\texttt{UNI}$ at approximately $\mathrm{R}_{\mathrm{a}}\approx 1.48$, giving rise to a second branch of heteroclinic-like POs HTPOT1b (see Fig. 9a ). We have observed that the two branches, HTPOT1a and HTPOT1b coexist in a range of activity numbers. Particularly, they coexist at $\mathrm{R}_{\mathrm{a}}\approx 1.5$.
(a) A member of the second branch of heteroclinic-like periodic orbits, HTPOT1b, which is $\sigma_{y}$-symmetric, shown close to the second heteroclinic bifurcation at $Ra\approx 1.5$, together with UNI and $\sigma_{y}$UNI, indicated by red and blue stars, respectively. Also shown are three velocity snapshots along the orbit: two exhibiting opposite unidirectional flows and one with a vortical flow state.
(b) Confirmation of the logarithmic scaling of period near the second heteroclinic bifurcation. Here, $A\approx 58.084$, $B\approx-18.039$, and $\mathrm{R}_{\mathrm{a_{c}}}\approx 1.48$.
Figure 9: A second branch of heteroclinic-like periodic orbits, HTPOT1b, and associated logarithmic scaling near the second heteroclinic bifurcation point.
III.4.5 Secondary Bifurcations Associated with Higher Unstable Eigenvalues of UNI
Our analysis so far has been limited to bifurcations from the primary unstable eigenspace of UNI. As the activity number is increased to $\mathrm{R}_{\mathrm{a}}=4.5$, a total of seven pairs of complex eigenvalues eventually cross the imaginary axis at various intermediate activity values. The corresponding secondary bifurcations give rise to additional ECSs, each embedded within a distinct symmetry subspace. While a full analysis is out of scope, it is crucial to identify how the symmetries of these other eigenspaces shape the bifurcation structure and the global dynamics deep into the turbulent regime.
Each crossing corresponds to a Hopf bifurcation, and the symmetries of each unstable 2D eigenspace are listed as isotropy groups in Table 3 in the Appendix. These symmetries constrain the nature of solutions that can emerge from bifurcations. Although it is not practical to trace every secondary bifurcation, we observe that RPOs with the corresponding symmetry commonly emerge, often via bifurcation scenarios similar to the SNIPER bifurcation analyzed in detail for the first eigenvalue pair.
For instance, the second unstable eigenvalue pair leads to a bifurcation scenario closely resembling that of the first pair, resulting in a branch of traveling waves with isotropy group $\Sigma_{p_{2(\text{U})}}$ (Fig. 10). This branch, which looks like two copies of the $TW$ (from the first pair) stacked horizontally, undergoes a saddle-node bifurcation and subsequently a SNIPER bifurcation. The result is a branch of homoclinic-like RPOs that we call HRPOT2a. However, in contrast to the first eigenvalue pair, no subsequent heteroclinic connections between UNI and $\sigma_{y}$UNI were observed upon increasing the activity.
The branches arising from bifurcations associated with the remaining eigenvalues populate the phase space with solutions that inherit the corresponding underlying symmetries. In particular, the second and third unstable eigenspaces have the symmetry $\tau_{x}(L/2)$, and therefore give rise to solutions confined to the invariant subspace defined by this transformation. As we shall see in the following sections, this subspace plays a crucial role in shaping the dynamics of reversals in the turbulent regime.
Figure 10: A snapshot of the nematic director and velocity fields of a traveling wave solution with isotropy group $\Sigma_{p_{2(\text{U})}}$, computed at $\mathrm{R}_{\mathrm{a}}=1.0$. This solution originates from UNI via an equivariant Hopf bifurcation at $\mathrm{R}_{\mathrm{a}}\approx 0.97$, associated with the second complex-conjugate pair of eigenvalues crossing the imaginary axis.
III.5 Secondary bifurcations from LAN
In this subsection, we carry out an equivariant bifurcation analysis of the secondary instabilities emerging from the LAN state in MFU A. Our main aim is to clarify the origin of dynamically relevant vortex–lattice–like ECSs. Oscillatory vortex states have been reported in several experimental and numerical studies [61, 34, 9, 26, 27]; in particular, Wagner et al. found vortex patterns whose characteristic length scale adjusts to the channel aspect ratio. A subset of these, termed ‘dancing disclinations’ [9, 68, 69, 70], exhibit +1/2+1/2 defects that form a space–time braid, and appear as POs and RPOs [26, 27]. Here we show that the relevant vortex–lattice solutions are not connected to the previously identified secondary bifurcations of the UNI equilibrium, and instead most likely arise from secondary bifurcations of the LAN equilibrium.
III.5.1 One-vortex equilibria
We consider the isotropy subgroup of LAN, $\Gamma_{LAN}=\mathbb{Z}_{2}(\sigma_{x}\sigma_{y})\rtimes SO(2)\cong O(2)$, as the base group for the equivariant analysis of bifurcations from LAN. The LAN solution first loses stability when a pair of real eigenvalues simultaneously crosses the imaginary axis at $\mathrm{R}_{\mathrm{a}}\approx 0.85$. Since the eigenvalues are real, this is a steady bifurcation, and the two-dimensional center subspace is taken as $\mathbb{R}^{2}$. In this center subspace, we can choose $v_{1}$ and $v_{2}$ such that the reflection $\sigma_{x}\sigma_{y}$ is represented by
$\sigma_{x}\sigma_{y}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. (9)
This is a reflection about the $x$-axis, which is also its fixed-point subspace. The translations act as the rotation matrices,
$R(\phi)=\begin{pmatrix}\cos\phi&-\sin\phi\\ \sin\phi&\cos\phi\end{pmatrix},\quad\phi=\frac{2\pi\ell}{L},\quad\ell\in[0,L),\quad\phi\in[0,2\pi).$ (10)
Every subgroup generated by the conjugate reflection $\Sigma_{\phi}=R(\phi)\sigma_{x}\sigma_{y}R(-\phi)$ is axial, and has the line (passing through origin) with slope $\phi$ as its fixed point subspace. The EBL predicts an infinite number of (pitchfork) branches of solutions bifurcating from LAN, one for each $\phi$, and possessing the corresponding $\Sigma_{\phi}$ symmetry.
Indeed, numerically we found a branch of steady one-vortex equilibrium solution (see Fig. 11a), invariant under the reflection $\Sigma_{0}=\sigma_{x}\sigma_{y}$, emerging from the LAN equilibrium. The other branches of solutions with $\Sigma_{\phi>0}$ symmetry are simply translations of this one-vortex solution. These solutions form a group orbit of equilibria, i.e., a relative equilibrium.
(a) The nematic director and velocity field of a one-vortex equilibrium solution bifurcated from the LAN at $\mathrm{R}_{\mathrm{a}}=0.85$. Its set of spatial translations constitutes a relative equilibrium.
(b) Two snapshots from a time simulation at $\mathrm{R}_{\mathrm{a}}=0.851$ of the stable one-vortex wave traveling to the left, separated by an interval of $t=50\tau$. This wave branch emerges via a steady bifurcation (of the corresponding normal vector field) from the one-vortex relative equilibrium.
Figure 11: One-vortex solutions: relative equilibria and traveling wave originating from the LAN solution in MFU A.
III.5.2 One-vortex traveling waves via symmetry-breaking bifurcation
The steady vortex solutions are stable for a small range of activity, $0.85\lessapprox\mathrm{R}_{\mathrm{a}}\lessapprox 0.851$. Like in the case of TW discussed earlier, the bifurcation occurring at $\mathrm{R}_{\mathrm{a}}\approx 0.851$ from the group orbit of the one vortex solution requires analyzing the vector field normal to the group orbit.
We denote any one-vortex equilibrium by $X_{0}$, and consider the corresponding kernel of Jacobian normal vector field, $d_{X}H(X_{0},0)$, which is one-dimensional and associated with a real eigenvalue in this case. The isotropy group acts nontrivially on this subspace, with the action of $\sigma_{x}\sigma_{y}$ given by $\mathbf{-1}$ [32, 63]. To apply the Equivariant Branching Lemma (EBL), we identify the axial group $\Sigma=\{I\}$, and predict a branch of solution with no symmetry. To find the group orbit of this solution, we apply Theorem 3 with $\Sigma=\{I\}$ with $\Gamma=O(2)$, the base group for LAN. This yields $N(\{I\})/\{I\}\cong O(2)$, which implies that the group orbit of this bifurcating solution is rank 1, and hence a traveling (rotating) wave, consistent with the general analysis for $O(2)$-equivariant systems [63].
(a) Family of periodic solutions bifurcating from the one-vortex steady state via a symmetry-preserving Hopf bifurcation, with the arrow indicating the parent steady solution. The colormap indicates activity. These periodic orbits grow in size as the activity is increased, and eventually collide with LAN and disappear via a homoclinic bifurcation.
[…]
Figure 12: One-vortex standing wave POs bifurcating from the steady state and terminating at a homoclinic bifurcation with LAN.
[…]
To the best of our knowledge, stable and coherent vortical traveling wave solutions have not been previously reported in active nematics. While active nematics are known to exhibit complex turbulent states characterized by transient vortices, the existence of stable, isolated traveling vortex waves represents a novel finding in this context.
III.5.3 One-vortex standing waves via symmetry-preserving bifurcation
The one-vortex TWs are solutions with no spatial symmetry. However, the previously reported dancing disclination solutions which are vortex dominated POs usually have the $\sigma_{x}\sigma_{y}$ symmetry [27]. To understand their origin, we look for symmetry-preserving bifurcations of the one-vortex equilibrium. We find that this equilibrium undergoes a second bifurcation at $\mathrm{R}_{\mathrm{a}}\approx 0.9$, when a complex pair of eigenvalues crosses the imaginary axis. The symmetry group $\Gamma=\mathbb{Z}_{2}(\sigma_{x}\sigma_{y})$ again acts trivially on the center subspace of the vector field $H$ normal to the group orbit.
Proposition 1 yields that $H$ undergoes a Hopf bifurcation to a unique time-periodic solution. To obtain the group orbit of the bifurcating solution, we apply Theorem 3 with $\Sigma_{X_{0}}=\mathbb{Z}_{2}(\sigma_{x}\sigma_{y})$ being the isotropy subgroup of the one-vortex solution (see Remark following Theorem 3), and $\Gamma=\Gamma_{LAN}$, to get $k=rank\left(\dfrac{\mathbb{Z}_{2}\times\mathbb{Z}_{2}}{\mathbb{Z}_{2}}\right)+1=rank(\mathbb{Z}_{2})+1=1$. This rules out RPOs (which have dimension 2), and implies that the bifurcating solutions of the full vector field are either standing waves (POs) or traveling waves (POs). An isotropy subgroup calculation similar to that performed for Hopf bifurcation from UNI shows that these solutions are in fact standing waves. These periodic solutions lie in the subspace defined by $\langle u\rangle=0$ and persist over a finite range of activity numbers. However, we find that these POs, while having the same symmetry as the dancing disclination solutions, do not have the characteristic braiding of $\dfrac{1}{2}$ defects of the latter [70].
Refer to captionRefer to caption(a) DPOT1a, Ra = 1.50(b) DPOT1b, Ra = 1.50
Figure 13: Two dancing-disclination-like POs (DPOs) that co-exist at $\mathrm{R_{a}}=1.5$, projected onto the reduced phase space ($\langle Q_{11}\rangle$$\langle v^{2}\rangle$). Both solutions lie in the $\langle u\rangle=0$ subspace (i.e., not net streamwise flow) due to the $\sigma_{x}\sigma_{y}$ symmetry.
III.5.4 Homoclinic bifurcation of the one-vortex POs and the emergence of dancing disclination solutions
As the activity is increased further, the size of one-vortex POs grows; the PO family eventually collides with the LAN and disappears via a homoclinic bifurcation at $\mathrm{R}_{\mathrm{a}}\approx 0.944$, as shown in Fig. 12a. The logarithmic divergence of the period near the homoclinic bifurcation confirms this observation (see Fig. 12b).
On the other side of the homoclinic bifurcation, we find a branch of one-vortex unstable POs with symmetry $\sigma_{x}\sigma_{y}$ around $\mathrm{R}_{\mathrm{a}}\approx 0.96$. This family, which we call DPOT1a, consists of exactly the dancing disclination type POs with the required braiding characteristics. While we currently do not understand the exact mechanism that connects the homoclinic bifurcation described above with the emergence of this branch, we hypothesize that the two are related due to their shared symmetry. Similarly, there exists another branch of periodic orbits invariant under the same symmetry, denoted DPOT1b, which also exhibits dancing defects. This branch of POs can be continued down to activity as low as $\mathrm{R}_{\mathrm{a}}=0.74$. A detailed analysis of the origin of these two branches of solutions is left for future work.
III.5.5 Secondary Bifurcations Associated with Higher Unstable Eigenvalues of LAN
So far, our analysis has been limited to bifurcations associated with the first unstable eigenspace of LAN. As the activity number is varied from $\mathrm{R}_{\mathrm{a}}\approx 0.85$ to $\mathrm{R}_{\mathrm{a}}\approx 4.5$, a total of eight (equal) pairs of real eigenvalues eventually cross the imaginary asix. The isotropy subgroups of these eight unstable 2D eigenspaces are listed in Table 4 in the Appendix. While the study of all bifurcations is outside the scope of this work, we briefly discuss the bifurcations from the second unstable eigenspace, since these solutions play an important role in the turbulent regime we analyze later in this paper.
This second pair of unstable eigenvectors possesses the $\tau_{x}(L/2)$ symmetry. The sequence of bifurcations is analogous to those of the first pair, and the bifurcating solutions possess the extra $\tau_{x}(L/2)$ symmetry. The steady solution is a two-vortex equilibrium that looks like two copies of the one-vortex equilibrium stacked along the $x$-axis. Additionally, there exists a branch of periodic orbits, denoted as POT2a (see Fig. 14), with symmetry $\sigma_{x}\sigma_{y}\tau_{x}(L/2)$ that cycles between a LAN-like and two-vortex like state. We anticipate that solutions associated with higher modes could bifurcate from LAN through similar mechanisms.
Figure 14: Visualization of POT2a at $\mathrm{R_{a}}=2.25$. (a) A snapshot of the nematic director and velocity fields, where its $\sigma_{x}\sigma_{y}\tau_{x}(L/2)$ symmetry is apparent. (b) Projection onto the reduced phase space: ($\langle Q_{11}\rangle$, $\langle v^{2}\rangle$). This PO lies in the $\langle u\rangle=0$ subspace (i.e., not net streamwise flow) due to the $\sigma_{x}\sigma_{y}$ symmetry.
IV Transient Flow Reversals in Preturbulent Flow
In section III, we uncovered the origin of several key ECSs that emerge from, or are related to, bifurcations of the UNI and LAN equilibria. We now discuss how these ECSs and their invariant manifolds facilitate robust, transient flow reversals in the preturbulent regime, where reversals are characterized by switching between near-unidirectional states located in the regions $\langle u\rangle<0$ (left) and $\langle u\rangle>0$ (right), respectively. By transient, we mean that each trajectory switches direction a finite number of times before entering the basin of attraction of a regular attractor (i.e. a stable ECS) or a chaotic attractor. By robust, we mean that the reversal mechanism persists for an open interval of activity values.
IV.1 Transient Reversal Dynamics at Low Activity: Organizing ECSs and Heteroclinic Connections
We focus on trajectories that begin near UNI or its symmetry counterpart $\sigma_{y}$UNI. When these ECSs are stable, all nearly uniaxial initial conditions decay to them. When these ECSs lose stability, the homoclinic-like RPOs (HRPOT1a and $\sigma_{y}$HRPOT1a) become the attractors instead, up to $\mathrm{R}_{\mathrm{a}}\approx 1.11$. Beyond this activity, and before the heteroclinic bifurcation that occurs at $\mathrm{R}_{\mathrm{a}}\approx 1.38$, the two RPOs are unstable. The unstable manifolds of these RPOs and the heteroclinic connections between these ECSs and the DPOs shape the flow between nearly uniaxial states.
[…]
(a) Mechanism 1: Flow reversal with no DPO mediator. The black dashed trajectory, initialized near UNI along its unstable manifold, successively follows the heteroclinic connections UNI $\rightarrow$ HRPOT1a $\rightarrow$ $\sigma_{y}$HRPOT1a and eventually settles into the localized chaotic attractor.
(b) Mechanism 2: Flow reversal using the DPOT1a as mediator. The black dashed trajectory follows the heteroclinic connections UNI $\rightarrow$ HRPOT1a $\rightarrow$ DPOT1a and settles into the localized chaotic attractor.
Figure 15: Two main mechanisms for flow reversals observed at $\mathrm{R}_{\mathrm{a}}=1.24$ by perturbing UNI.
Increasing the activity further reveals additional reversal mechanisms. At $\mathrm{R}_{\mathrm{a}}=1.5$, for instance, we observe that eight fundamental ECSs organize the reversal dynamics between left- and right-flowing states. These include the two heteroclinic-like HTPOT1a and HTPOT1b, the homoclinic-like HRPOT2a, the DPO1a and DPO1b associated with LAN, and three additional RPOs. These new RPOs, which we call DRPOT1a,b,c, likely emerge from bifurcations of the DPO branches. These are the drifting versions of dancing disclination solutions, as first reported in [26]. Interestingly, at this activity number, no localized chaotic attractor was found. Instead, the three DRPOs occupy approximately the same region of phase space that was previously occupied by the chaotic attractor, and DRPOT1c is the only attractor in this region.
By perturbing UNI along its two pairs of unstable directions at $\mathrm{R}_{\mathrm{a}}=1.5$, we obtain trajectories that initially shadow either HTPOT1b or HRPOT2a. In the first case, the initial transient consists of repeated reversals between the two nearly uniaxial flow states, due to the heteroclinic nature of HTPOT1b. Eventually, the trajectory diverges from HTPO1b and either moves directly to the region populated by the three DRPOs, or passes near the DPOT1a before reaching a symmetry instance of the DRPO family as shown in Fig. 16a. In the second case, after transiently shadowing HRPOT2a, a trajectory can undergo a reversal event by a direct heteroclinic connection to the DRPOT1a copy on the opposite side, or via an intermediate heteroclinic connection to DPOT1a. All trajectories at $\mathrm{R}_{\mathrm{a}}=1.5$ eventually end up in the attractor DRPOT1c or its $\sigma_{y}$ copy.
[…]
(b) Mechanism 1 using HRPOT2a. The black dashed trajectory follows the heteroclinic sequence UNI $\rightarrow$ HRPOT2a $\rightarrow$ $_{T1a}\rightarrow DRPO$T1b $\rightarrow$ DRPOT1c.
Figure 16: Main mechanisms of flow reversals at $\mathrm{R}_{\mathrm{a}}=1.5$ observed by perturbing UNI along its four dimensional unstable subspace. (a) A mechanism that uses the HTPOT1b with DPOT1a as the mediator. (b) A mechanism using HRPOT2a without DPOs as mediators.
In summary, for low activity numbers, three main types of reversal mechanisms are identified: the first involves heteroclinic connections originating from HRPOs or HTPOs to the relevant attractor; the second mechanism is similar but with DPOs serving as mediators; and the third arises when trajectories shadow either of the two HTPOs themselves. When both the HTPOs exist, we observe that HTPOT1b is more dominant of the two in influencing trajectories originating from perturbations of UNI. Animations of representative reversal trajectories and their shadowed ECS solutions are provided in the Supplemental Material.
IV.2 Transient Flow Reversals at Higher Activity and the Path Toward Turbulence
While tracking bifurcations all the way into the turbulent regime is impractical, we make a few observations regarding the role of the various symmetries in shaping the phase space structure as the activity number is increased.
We have seen earlier that solutions with symmetries conjugated to $\sigma_{x}\sigma_{y}$, and branches bifurcating from them, provide mechanisms for reversal. By following the DPOT1a branch into the higher activity region, we observe that at $\mathrm{R}_{\mathrm{a}}\approx 2.38$, this branch becomes stable. Then, at $\mathrm{R}_{\mathrm{a}}\approx 3.29$, it undergoes a period-doubling bifurcation that breaks the $\sigma_{x}\sigma_{y}$ symmetry, giving rise to a stable preperiodic orbit (see Fig. 17a). The resulting solution represents, by itself, a stable mechanism for periodic flow reversals about $\langle u\rangle=0$. As the activity number reaches $\mathrm{R}_{\mathrm{a}}=4.0$, new ECSs emerge from a sequence of bifurcations related not only to DPOT1a but also to nearby solutions, which increasingly populate this region of phase space in the form of a localized chaotic attractor shown in Fig. 17b. This sequence preludes the birth of the set $\mathcal{Y}$ at $\mathrm{R}_{\mathrm{a}}=4.5$ (see Fig. 19), within which solutions related to DPOT1a and their manifolds continue to play an important role in mediating reversals.
(a) The unstable DPOT1a at $\mathrm{R}_{\mathrm{a}}=3.5$ is shown in green. An attracting periodic orbit that emerges from DPOT1a at $\mathrm{R}_{\mathrm{a}}\approx 3.29$ via a period doubling bifurcation is shown in black. Notice that this new PO not in the plane $\langle u\rangle=0$, i.e., it breaks the $\sigma_{x}\sigma_{y}$ symmetry. Also note the extent of the $\langle u\rangle$ axis is much smaller than in earlier plots showing the HRPOs and HTPOs earlier.
(b) A trajectory attracted to a region populated by ECSs at $\mathrm{R}_{\mathrm{a}}=4.0$, serving as a precursor to the chaotic set $\mathcal{Y}$ at $\mathrm{R}_{\mathrm{a}}=4.5$.
Figure 17: Phase space visualizations illustrating the evolution from a periodic orbit to the onset of the chaotic set $\mathcal{Y}$.
The other set of ECSs that remain relevant at higher activity are those with $\tau_{x}(L/2)$ symmetry, emerging from bifurcations of the UNI and LAN branches. In particular, at $\mathrm{R}_{\mathrm{a}}=3.75$, there exists a periodic orbit that is stable within the invariant subspace under this symmetry and attracts perturbations of UNI along the second and third pairs of unstable eigenvectors. It seems to be related to a secondary bifurcation of POT2a that breaks the $\sigma_{x}\sigma_{y}$ symmetry. Interestingly, at $\mathrm{R}_{\mathrm{a}}\approx 4.0$, no ECSs within this subspace are stable, and a new chaotic set emerges. We denote this set by $\mathcal{X}$, as shown in Fig. 19. At this stage, trajectories initialized from UNI and perturbed within the four-dimensional subspace composed of the 2D second and third unstable eigenspaces can cross into the other region of phase space, near $\sigma_{y}$UNI, and switch recurrently between these two states.
Figure 18: At $\mathrm{R}_{\mathrm{a}}=3.75$, the black dashed trajectory perturbed from UNI settles into a periodic orbit (magenta) that is stable within the $\tau_{x}(L/2)$ invariant subspace. This periodic orbit seems to have originated from secondary bifurcations of POT2a. The periodic orbit POT2a is shown in green.
The chaotic sets $\mathcal{X}$ and $\mathcal{Y}$ exhibit reversals between regions where the flow velocity has a mean value $\langle u\rangle$ that persistently oscillates about positive and negative values, respectively. In the following section, we analyze these attractors to elucidate the mechanisms responsible for flow reversals in the MUF at $\mathrm{R}_{\mathrm{a}}=4.5$.
V Flow Reversal Mechanisms in Turbulence: Two-Way Persistent Reversals
(a) Chaotic set $\mathcal{X}$
(b) Chaotic set $\mathcal{Y}$
(c) Quasiperiodic set $\mathcal{Z}$
Figure 19: All perturbations of UNI along its unstable manifold in the MFU A end up in one of the three sets $\mathcal{X}$, $\mathcal{Y}$ and $\mathcal{Z}$. The top row shows the chaotic sets $\mathcal{X}$ and $\mathcal{Y}$. The bottom row shows the quasiperiodic set $\mathcal{Z}$.
The results presented for the preturbulent regime, where we have shown that the dynamical trajectories are shaped by the ECSs and their invariant manifolds, motivate us to investigate the nature of the reversal mechanisms within the turbulent regime. In the original full-channel configuration (length $\tilde{L}$), however, Wagner et al. [27] explored the turbulent regime at $\mathrm{R}_{\mathrm{a}}=4.5$, reporting no evidence of shadowing among the 210 distinct ECSs and their discrete-symmetry counterparts, leaving the role of ECSs in the dynamics still unclear. In contrast, our use of smaller domains as MFUs allows us to identify shadowing events in a simplified system with fewer solutions and reduced complexity. These detected shadowing events highlight the pathways through which chaotic trajectories undergo reversals. In this section, we consider two MFUs: the MFU A which has been analyzed throughout the previous sections with length $L_{A}=\tilde{L}/3$, and its halved version MFU B with $L_{B}=\tilde{L}/6$. We explore the following questions in the following:
  • Q1: Where do typical trajectories end up at $\mathrm{R}_{a}=4.5$ in the two MFUs ?
  • Q2: Do the trajectories that end up in one of the two chaotic sets $\mathcal{X}$ or $\mathcal{Y}$ shadow the ECSs embedded in those sets at $R_{a}=4.5$?
  • Q3: What are the symmetries and possible origins of ECSs embedded in the chaotic sets ?
The UNI in the MFU A at $\mathrm{R}_{\mathrm{a}}=4.5$ has seven complex pairs of unstable eigenvalues. Perturbations of the UNI within the union of eigenspaces corresponding to the first, fourth and fifth pairs evolve towards the chaotic set $\mathcal{Y}$, while perturbations within the union of eigenspaces of the second and third pairs evolve towards the chaotic set $\mathcal{X}$. In the remaining two directions, perturbations lead to a stable quasiperiodic orbit in the MFU, denoted by $\mathcal{Z}$, see Fig. 19.
We find that an arbitrary perturbation (i.e., with no symmetry) always ends up in $\mathcal{Y}$, indicating that in MFU A, $\mathcal{Y}$ is the only almost-global attractor while the other two sets $\mathcal{X}$ and $\mathcal{Z}$ are saddles. Motivated by the apparent $\tau_{x}(L_{A}/2)$ symmetry of $\mathcal{X}$ itself, we verify that in fact, all initial conditions with $\tau_{x}(L/2)$ symmetry end up in $\mathcal{X}$ in MFU A. Hence, we also analyze dynamics in the smaller MFU B, where $\mathcal{Y}$ is eliminated due to lack of two-fold translation symmetry, and we find that the halved-version of $\mathcal{X}$ is the only almost-global attractor that attracts arbitrary initial conditions almost surely. The quasiperiodic $\mathcal{Z}$, while still present in MFU B, again only attracts a measure zero set of trajectories that have the additional translation symmetry. Table 1 summarizes this discussion and answers Q1 posed above. Now we turn to answering Q2 and Q3.
V.1 Establishing shadowing within the chaotic sets $\mathcal{X}$ and $\mathcal{Y}$
To establish shadowing, we perform long-time simulations ($t=30000\tau$) in MFU A within each of the two chaotic sets and focus on the final segment, $29000\tau,30000\tau]$, where transient effects have long decayed. Both chaotic sets exhibit persistent flow reversals between two nonzero mean flow states, as evidenced by their reduced phase space projections and the time series of representative trajectories over this interval. We analyze shadowing events involving a dominating set of 75 ECSs in MFU A. Within this set, we identify four symmetry-based families: 40 ECSs with no additional translation symmetry (T1), 32 ECSs invariant under $\tau_{x}(L_{A}/2)$ (T2), one ECS invariant under $\tau_{x}(L_{A}/3)$ (T3), and two ECSs invariant under $\tau_{x}(L_{A}/4)$ (T4). Each family includes the corresponding symmetry-related counterparts, and we investigate how these families contribute to organizing the dynamics within the chaotic attractors.
For the plots in this section, we color the ECS based on the spatial average of the streamwise velocity $\langle u\rangle$ as follows:
  • ECSs confined to $\langle u\rangle>0$ (red),
  • ECSs confined to $\langle u\rangle<0$ (blue),
  • ECSs crossing the subspace $\langle u\rangle=0$ (green).
To establish shadowing and identify shadowed ECSs, we follow a two step approach [72]. In the first step, for the final $1000\tau$ window within these attractors, we analyze the distances between a typical turbulent trajectory and every ECS (including translation-symmetry-related copies) using a continuous-symmetry-reduced distance [27],
$d(X_{1}(t),X_{E})=\min_{\begin{subarray}{c}0\leq\ell\leq L\\ 0\leq s\leq T\end{subarray}}\left\|\tau_{x}(\ell)X_{1}(t)-X_{E}(s)\right\|_{2}$ (11)
where $X_{1}(t)$ is the turbulent trajectory and $X_{E}(t)$ is the trajectory of the corresponding ECS. At each time step, we identify the ECS with the minimum distance to the turbulent trajectory. The turbulent trajectory should exhibit dynamics similar to those of the ECS during the shadowing event [31, 73]. In the second step, we confirm shadowing by comparing the dynamical evolutions with the trajectory for the ECSs selected as closest to the trajectory, as we discuss next.
Length $\mathcal{X}$ $\mathcal{Y}$ $\mathcal{Z}$
MFU A $L_{A}=\tilde{L}/3$ Chaotic saddle Chaotic attractor Quasiperiodic saddle
$\tau_{x}(L_{A}/2)$ symmetric No symmetry $\tau_{x}(L_{A}/4)$ symmetric
MFU B $L_{B}=\tilde{L}/6$ Chaotic attractor Doesn’t exist Quasiperiodic saddle
No symmetry $\tau_{x}(L_{B}/2)$ symmetric
Table 1: Properties of the three invariant sets in the two MFUs. For MFU B, the entries refer to the corresponding sets truncated to the reduced domain.
V.2 Chaotic Set $\mathcal{X}$
V.2.1 Dominant ECSs Within $\mathcal{X}$
We consider a typical trajectory $X(t)$ lying in $\mathcal{X}$ in MFU A over the interval $29000\tau,30000\tau]$. This segment of the trajectory exhibits eleven flow reversals. The mean streamwise velocity $\langle u\rangle$ fluctuates around zero, with extended excursions near $\langle u\rangle=-2$ (net negative flow) and $\langle u\rangle=2$ (net positive flow). To investigate the underlying structure, we record the closest ECS to the trajectory every $0.4\tau$, yielding 2501 comparison points. These ECSs are color-coded as discussed above. As shown in Fig. 20, when the trajectory transitions between red and blue regions, it typically passes near a green ECS.
(a) Time series of $\langle u\rangle$ along the turbulent trajectory $X(t)$ over the interval $29000\tau\leq t\leq 30000\tau$. Various shadowing events are also marked, colored by the region occupied by the corresponding shadowed ECS. The encounters with ECS that lie completely in the half-space ($\langle u\rangle>0$) are in red, with ECS that lie completely in the half-space ($\langle u\rangle<0$) are in blue, and with ECS that lie on both sides of the hyperplane ($\langle u\rangle=0$) are shown in green.
(b) Reduced phase space projection showing the segment turbulent trajectory $X(t)$ in the interval $29000\tau\leq t\leq 30000\tau$ (black) and the six main ECS groups. The color scheme for the ECSs is the same as in (a). The orange circle indicates the beginning of the segment ($t=29000\tau$) and the magenta square indicates the end of the segment ($t=30000\tau$).
(c) Reduced phase space portrait showing the six main ECS groups shadowed by the turbulent trajectory, along with their symmetry-related copies under $\sigma_{y}$.
Figure 20: Flow reversals and the relationship between the turbulent trajectory and the shadowed ECSs.
Across the 2,501 points, 40 of the 75 ECS groups appear as the closest state at least once. To isolate the most dynamically significant structures, we focus on the ten most frequently encountered ECS groups. Together, these ten groups account for approximately 62% of all occurrences (see Table 6 in Appendix). As anticipated, each of these ten groups consists of ECSs that are invariant under $\tau_{x}(L_{A}/2)$ symmetry.
Figure 21: Symmetry-reduced distance plot showing the closest encounters between the chaotic trajectory $X(t)$ and the ten relevant ECS and their symmetry counterparts. The ECS list on the left follows the same color scheme as in Fig. 20(a). The plot shows a clear pattern of repeated transitions (i.e., reversals) between the ‘left flowing’ (blue) and ’right flowing’ (red) ECSs, via the intermediate (green) ECSs.
These ten ECS groups serve as a reduced representation that captures the essential skeleton of the attractor $\mathcal{X}$’s dynamics. Six of these groups belong to the green category and are natural candidates for mediating transitions between the red and blue regions, i.e., for enabling reversals. Figure 21 shows the distance plot between the trajectory and these ECS groups, highlighting the paths of reversals. In particular, the four ECS groups from the red/blue categories, along with those from the green category, are closely approximated by the trajectory. For example, in the interval $29000\tau,29200\tau]$, during the first reversal of the series (see also Fig. 20a), the distance plot shows that the trajectory first approaches the blue ECS group, with the mean velocity oscillating around $\langle u\rangle\approx-2$. It then nears some ECSs of the green group, where $\langle u\rangle\approx 0$, before finally transitioning to a state where the mean velocity oscillates around $\langle u\rangle\approx 2$ and closely follows the red ECS group.
V.2.2 Shadowing in $\mathcal{X}$
Among the above chosen subset of ECSs, we identify the closest ECS at each time step within the interval $29000\tau,30000\tau]$, and assess whether these close encounters represent genuine shadowing events. To make this determination, we compare the dynamics of the trajectory with those of the ECSs. Specifically, we utilize the time parameter $s$ associated with each ECS, along with the streamwise translation $\ell$, which is applied to obtain the closest translated version of the ECS as defined in Equation 11. A shadowing event is characterized by synchronized evolution between the ECS and the turbulent trajectory. Thus, during a shadowing event, we expect $\dfrac{ds}{dt}\approx 1$ and $\dfrac{d\ell}{dt}\approx 0$, indicating that the ECS evolves at nearly the same rate as the trajectory and requires minimal translation to remain aligned with the turbulent trajectory.
Since turbulent trajectories inevitably depart from the ECS along its unstable manifold, we adopt a relaxed criterion to identify approximate shadowing. Specifically, the criterion is considered satisfied when
$0.5\leq\frac{ds}{dt}\leq 1.5,$ (12)
as illustrated in Fig. 22. In such cases, we typically observe
$\frac{d\ell}{dt}<1,$ (13)
which corresponds to a shift of less than 4% of the channel length. Time derivatives are computed using central differences for interior points and forward/backward differences at endpoints.
For relative periodic orbit ECSs, we additionally account for the discrete jumps in $s$ and $\ell$ at the orbit period $T$:
$\tau_{x}(\ell_{0})X_{E}(s=T)=X_{E}(s=0),\quad\text{so that}\quad s\rightarrow s+T,\;\;\ell\rightarrow\ell+\ell_{0},$ (14)
which preserves continuity in the symmetry-reduced representation of the RPO and allows consistent computation of the derivatives.
Figure 22: The dynamic shadowing criterion evaluated along the trajectory $X(t)$ relative to the closest ECS $X_{E}(s)$ at discrete times (with $\delta t=0.4\tau$) within the interval $29000\tau,30000\tau]$. The plot marks instances where the dynamical similarity measure $\dfrac{ds}{dt}$ lies in the interval $[0.5,1.5]$, thereby satisfying the criterion. The closer its value is to $1$, the stronger the shadowing event.
Figure 23: Three snapshots for three different confirmed shadowing events of trajectory $X(t)$. The first row of each event shows sequential snapshots of the turbulent trajectory, while the second row shows sequential snapshots (at the same times) of the shadowed ECS. Notice the translation $l$ is almost constant during each event, which implies that the same translated copy an ECS is being shadowed throughout the time interval.
Shadowing is typically most evident when the distance between the turbulent trajectory and an ECS remains within a small range (typically between $0.07$ and $0.3$). In the interval $29099.2\tau,29104.4\tau]$ (see Fig. 23(a)), the ECS closest to $X(t)$ is $\mathrm{T}2/4$, with an average distance $\langle d\rangle\approx 0.247$ and average rates $\langle|\dfrac{ds}{dt}|\rangle\approx 1.055$ and $\langle|\dfrac{d\ell}{dt}|\rangle\approx 0.101$. During this interval, the trajectory reaches the ECS within a minimum distance of $d\approx 0.2$ at $t=29102\tau$, where $\dfrac{ds}{dt}\approx 1$ and $\dfrac{d\ell}{dt}\approx-0.005$, which are the optimal values of these indicators within this interval. Taken together, the consistently small distance, near-unity $\dfrac{ds}{dt}$, and small $\dfrac{d\ell}{dt}$ throughout this interval provide strong confirmation of a shadowing event.
Another clear shadowing event occurs in the interval $29799.2\tau,29806.8\tau]$ (see Fig. 23(b)). In this case, the closest ECS is $\mathrm{T}2/15$, with average distance and rates $\langle d\rangle\approx 0.126$, $\langle|\dfrac{ds}{dt}|\rangle\approx 0.989$, and $\langle|\dfrac{d\ell}{dt}|\rangle\approx 0.183$. At $t=29805.6\tau$, the trajectory has $d\approx 0.105$, $\frac{ds}{dt}\approx 1.025$, and $\frac{d\ell}{dt}\approx 0.017$.
When the shadowing criteria are satisfied at larger distances, the resemblance to the ECS is generally weaker. For example, around $t=29939.6\tau$ (see Fig. 23c), when the turbulent trajectory is in the red region, it does not approach the ECS $\mathrm{T}2/13$ closer than $0.453$, yet the temporal evolution remains similar, with $\frac{ds}{dt}\approx 0.813$ and $\frac{d\ell}{dt}\approx 0.391$. Although the resemblance to the ECS is not as strong as in cases with smaller distances, these example suggest that, even at relatively large distances, traces of shadowing can still be detected. This could indicate that the trajectory is closely shadowing another, unconsidered but dynamically related ECS associated with the considered one, or possibly shadowing an invariant manifold instead. Therefore, the analysis of shadowing reveals different nuances that must be taken into account due to the complexity of the phase space geometry.
V.2.3 Reversal Mechanisms and Heteroclinic connections
The shadowing events confirmed above hint at mechanisms similar to those discussed for the preturbulent regime in Section IV. Consider, for instance, the first clear reversal event in $X(t)$, occurring within the interval $29034\tau,29062.8\tau]$ (see Fig. 20a). The system transitions from a leftward to a rightward unidirectional flow, while passing close to a two-vortex like ECS. Figure 24 illustrates three representative snapshots capturing the key stages of this reversal event, along with their corresponding shadowed ECSs, highlighting how the trajectory temporarily aligns with these coherent structures. Figure 25 complements this view by showing the reduced phase space representation of the same event.
Figure 24: Top row: Velocity field snapshots during the first reversal event for a turbulent trajectory $X(t)$, lying in attractor $\mathcal{X}(t)$, during the interval $29034\tau,29062.8\tau]$ resembling the left UNI, two-vortex lattice and right UNI solutions, respectively. Bottom row: Velocity field snapshots of the ECSs shadowed during the reversal shown at the respective times.
To obtain more insight into the reversal dynamics, the network of heteroclinic connections among the ECSs needs to be studied. However, even in the reduced representation involving a set of 40 ECSs (modulo symmetries), the transition pathways between the blue, green, and red regions are too complicated to be explicitly traced. But some aspects of the phase space geometry become clearer when we focus on specific pairs of ECSs.
First, the green ECSs play a role analogous to that of the ‘intermediary’ DPOs in the preturbulent regime. These green ECSs are likely originated from solutions invariant under symmetries conjugate to $\sigma_{x}\sigma_{y}$ and $\tau_{x}(L_{A}/2)$ in the preturbulent regime. These ECSs in the turbulent regime have slightly broken $\sigma_{x}\sigma_{y}$ symmetry, but retain the two-fold $\tau_{x}(L/2)$ translational symmetry. Supporting evidence includes their similarity to POT2a, and the observation that when a green solution is shadowed, its $\sigma_{x}\sigma_{y}$ counterpart is also shadowed (see Fig. 21) simultaneously, implying that these solutions with broken $\sigma_{x}\sigma_{y}$ symmetry lie close to each other. Second, the blue and red ECSs play a role analogous to that of HRPOs, and form the structural backbone for generating left- and right-flowing, nearly uniaxial states. Their phase space structures suggests that these ECSs are most likely the result of symmetry-breaking bifurcations of the HRPOs. The blue/red ECS, which retain the homoclinic-like structure in phase space, have broken the shift-reflect symmetries of their preturbulent counterparts, but still possess the two-fold $\tau_{x}(L_{A}/2)$ translation symmetry. Finally, the heteroclinic connections between these groups are likely the mechanisms through which turbulence exhibits flow reversals.
Figure 25: Shadowing and reversals of Fig. 24 in the reduced phase space. A segment of the turbulent trajectory $X(t)$ is shown in black, while the shadowed left-flowing ECS $\sigma_{x}\sigma_{y}$T2/18 is shown in blue, the two-vortex ECS $\sigma_{x}\sigma_{y}$T2/4 is shown in green, and the right-flowing $\sigma_{x}\sigma_{y}$T2/1 is shown in red. The larger circle, rhombus, and square mark the locations of the snapshots (inset) along the turbulent trajectory, while the smaller circle, rhombus, and square indicate the respective locations along the shadowed ECSs. Additionally, the red and blue asterisks show the UNI and $\sigma_{y}$UNI, respectively. An animation of this reversal event is provided in the Supplemental Material.
Indeed, we observe that heteroclinic connections may exist between pairs of ECSs from different regions that are shadowed sequentially. For example, the red ECS $\sigma_{x}\sigma_{y}$T2/1 is shadowed during the interval $20067.20\tau,\,29073.6\tau]$, and immediately afterward, the green ECS $\sigma_{y}$T2/30 is shadowed during $29072.80\tau,\,29073.6\tau]$. This suggests a heteroclinic connection between these two ECSs. However, within the green region, the turbulent trajectory sequentially (and sometimes simultaneously) shadows several ECSs, including $\sigma_{y}$T2/30, $\sigma_{x}$T2/25, T2/13, $\sigma_{y}$T2/9, $\sigma_{x}$T2/30, $\sigma_{x}\sigma_{y}$T2/4, and $\sigma_{x}\sigma_{y}$T2/13. Consequently, identifying a heteroclinic connection to a target green ECS is challenging, as shadowing events at the red/green and blue/green boundaries involve multiple ECSs and are not sharply defined. Although we cannot conclusively verify a heteroclinic connection between $\sigma_{x}\sigma_{y}$T2/1 and $\sigma_{y}$T2/30, the turbulent trajectory shadows the unstable manifold of the former ECS and eventually approaches the latter. Throughout the subsequent exploration of the green region, the trajectory continues to follow the unstable manifold, as illustrated by the segment of $X(t)$ and a nearby trajectory in Fig. 26 within the interval $29064.0\tau,\,29094.4\tau]$.
Figure 26: Shadowing of the unstable manifold of a ECS by a turbulent trajectory in $\mathcal{X}$. The turbulent trajectory (black), shadows a trajectory (red dashed) on the unstable manifold of a ECS, $\sigma_{x}\sigma_{y}$T2/1, during $29064\tau,29094.4\tau]$. The ECS itself is shown in solid red. The later segment of the turbulent trajectory where it shadows a second ECS, $\sigma_{y}$T2/30, is shown in green. This latter ECS omitted from the plot for clarity. An animation of this shadowing event is provided in the Supplemental Material.
The previous example illustrates how the ECSs and their invariant manifolds shape the dynamics of transitions. To obtain evidence of connections between the green and blue/red ECSs, we perform our computations in MFU B. Recall that in MFU B, $\mathcal{X}$ is the only attractor. Additionally, the ECSs and connections that exist in MFU B and determine the attractor $\mathcal{X}$ dynamics are necessarily those that possess $\tau_{x}(L_{A}/2)$ symmetry in MFU A.
In this MFU B based calculation, we focus on the two most relevant blue/red ECSs and the three most relevant green ECSs (and their symmetry counterparts). In this smaller domain, the RPOs T2/1 and T2/15 each have two unstable directions, while the RPOs T2/9, T2/30, and T2/4 have three. We consider the dominant unstable direction to generate a sample of trajectories within each unstable manifold. Several of these trajectories pass nearby and shadow other ECSs, providing evidence of heteroclinic connections. Table 2 lists few of the pairs of ECSs connected by such trajectories, together with the corresponding minimum distances and shadowing time intervals. The Figs. 27a and 27b highlight two such examples.
Table 2: Examples of trajectories on the unstable manifolds of source ECSs that eventually shadow target ECSs. These trajectories connect the red/blue (left-flowing/right-flowing) group of ECSs to the green (intermediate) group of ECSs. Some of these trajectories appear to approach heteroclinic connections between the source and target pairs.
Source ECS Target ECS Type Min. distance Shadowing interval
T2/9 $\sigma_{x}\sigma_{y}$T2/1 green $\rightarrow$ red 0.043 $10.4\tau$
T2/9 $\sigma_{y}$T2/15 green $\rightarrow$ blue 0.053 $22.4\tau$
T2/30 $\sigma_{y}$T2/1 green $\rightarrow$ red 0.061 $12.8\tau$
T2/30 $\sigma_{x}\sigma_{y}$T2/15 green $\rightarrow$ blue 0.009 $19.6\tau$
T2/4 $\sigma_{x}\sigma_{y}$ T2/1 green $\rightarrow$ red 0.062 $12.4\tau$
T2/4 $\sigma_{y}$T2/15 green $\rightarrow$ blue 0.010 $22.4\tau$
T2/15 $\sigma_{y}$T2/9 red $\rightarrow$ green 0.106 $13.2\tau$
T2/15 $\sigma_{x}$T2/30 red $\rightarrow$ green 0.15 $13.6\tau$
T2/15 T2/4 red $\rightarrow$ green 0.115 $9.6\tau$
T2/1 $\sigma_{x}\sigma_{y}$ T2/9 blue $\rightarrow$ green 0.157 $14.8\tau$
T2/1 T2/30 blue $\rightarrow$ green 0.106 $14\tau$
T2/1 $\sigma_{x}$T2/4 blue $\rightarrow$ green 0.167 $16\tau$
Notice that the paths from the green to the red/blue ECSs are clearer than those from the red/blue to the green ECSs. This is related to the fact that the green region is populated by a larger number of unstable ECSs that interact with each other in a more complex manner. For example, the minimum distances between the trajectory on the unstable manifold of T2/30 that shadows T2/15 and the trajectory from T2/9 to T2/1 are on the order of $10^{-3}$ and $10^{-2}$, respectively (see Table 2), which is remarkable given the level of simplification used to identify them. In contrast, the trajectories originating from the unstable manifolds of T2/1 and T2/15 do not get sufficiently close to a heteroclinic trajectory to confirm a connection; nevertheless, they clearly illustrate the dynamic paths from the red/blue to the green regions by exhibiting a clear shadowing of the latter. These results demonstrate that the unstable manifold of the source ECS eventually shadows the target ECS. This finding confirms that the ECSs and their invariant manifolds constitute the scaffold underlying reversals in the turbulent set $X$.
Since $\mathcal{X}$ is the only attractor in MFU B, the above discussion implies that (a) arbitrary trajectories in MFU B ultimately converge to $\mathcal{X}$, and (b) trajectories on this attractor repeatedly shadow various ECSs and portions of their invariant manifolds.
(a) T2/9 $\rightarrow\sigma_{y}T2/15$
(b) T2/30 $\rightarrow\sigma_{x}\sigma_{y}T2/1$
Figure 27: Examples of trajectories on the unstable manifolds of intermediary (green) ECSs T2/30 and T2/9 that eventually shadow the left-flowing $\sigma_{y}T2/15$ (blue) and right-flowing $\sigma_{x}\sigma_{y}T2/1$ (red) ECSs, respectively. The minimum distance between a trajectory and the target ECS was found to be $0.043$. These results provide strong evidence for the existence of the corresponding heteroclinic connections.
V.3 Chaotic Set $\mathcal{Y}$
V.3.1 Approximations to the ECSs of a Typical Trajectory in $\mathcal{Y}$
We now consider a trajectory $Y(t)$ within the chaotic set $\mathcal{Y}$ over the interval $29000\tau,30000\tau]$ in MFU A. Several key differences are evident when compared to a typical trajectory $X(t)$ in $\mathcal{X}$. First, the oscillations in $Y(t)$ are shorter, and the maximum value of $|\langle u\rangle|$ is about 3, compared to about 4 for $X(t)$. This suggests a less pronounced reversal scenario, with the separation between the blue/red and green regions being less distinct than in $X(t)$. Second, while $Y(t)$ occasionally approximates blue/red ECSs when near $\langle u\rangle\approx 0$, this occurs less frequently than in $X(t)$. Thus, the overlap between the green ECSs and blue/red ECSs is more significant in $Y(t)$, preventing a clear distinction of ECS groups that mediate the reversals. Third, and most importantly, the minimum distances between the trajectory and the ECSs are larger for $Y(t)$ than for $X(t)$.
Figure 28: Time series of $\langle u\rangle$ along the turbulent trajectory $Y(t)$. Various shadowing events are also marked, colored by the region occupied by the corresponding shadowed ECS, using the same color scheme as in Fig. 20(a).
In general, the shadowing of the ECSs is less pronounced in $Y(t)$, as shown in Table 7 in the Appendix for the fifteen most frequent visited ECS groups. For example, the minimum distance from the trajectory to the most frequently visited ECS is 0.308, whereas the minimum distance to the most frequent visited ECSs for $X(t)$ is only 0.119. Notice that we also needed to include more ECSs to cover about $62\%$ of the closest approaches in $Y(t)$ than we did for $X(t)$.
The fact that the set of 75 ECSs is not exhaustive may explain why we do not observe a clearer picture of the dynamical reversals for $\mathcal{Y}$. Indeed, the distance plot in Fig. 29, for the fifteen most visited ECS groups, shows no clear pattern or contrast between regions, unlike the plot in Fig. 21, which may be attributed to the absence of ECSs that have not yet been computed. The attractor $\mathcal{Y}$ is not constrained by the $\tau_{x}(L/2)$ symmetry of $\mathcal{X}$. The additional complexity arises in part from this doubling of the effective MFU length. This is consistent with the principal component analysis (PCA) of the corresponding trajectories (see Table 8): for $X(t)$, nearly 85% of the variance is captured by just the first two principal components (PCs), and 90% is reached with four; in contrast, $Y(t)$ requires eight components to explain 90% of the variance, indicating that it explores a significantly higher-dimensional region of phase space.
Figure 29: Symmetry-reduced distance plot showing the closest approaches between the turbulent trajectory $Y(t)$ and 15 dynamically relevant ECSs (and their symmetry counterparts). Together, these ECSs account for more than 60% of all shadowing events. In contrast to the trajectory $X(t)$ (see Fig. 21) where the shadowing events were well separated in time, here the shadowing of the green ECSs and the red/blue ECSs appears to occur concurrently. The minimum distance attained is 0.24, roughly three times larger than the smallest distance observed for $X(t)$.
V.3.2 Shadowing in $\mathcal{Y}$
Although the reversal dynamics are not fully captured due to the reduced set of relevant ECSs and the inherent complexity of the chaotic set $\mathcal{Y}$, our analysis still partly uncovers the underlying deterministic skeleton. Figure 30 shows that shadowing events are less frequent than those of the $X$-trajectory. Nevertheless, some of these events exhibit dynamics that closely resemble the behavior of the corresponding ECSs, as illustrated in the examples presented in Fig. 31.
The first example occurs in the interval $29065.2\tau,29066.4\tau]$, where $Y(t)$ shadows $\sigma_{y}\mathrm{T1/2}$, with $\langle d\rangle\approx 0.443$, $\langle|ds/dt|\rangle\approx 0.921$, and $\langle|d\ell/dt|\rangle\approx 0.392$. In Fig. 31a, we show three consecutive snapshots from this interval, corresponding to the moment when the distance reaches its minimum value $d\approx 0.393$. The second example in Fig. 31 is a shadowing of $\sigma_{x}\mathrm{T1/3}$ within the interval $29602.8\tau,29603.6\tau]$. In this case, $\langle d\rangle\approx 0.304$, $\langle|ds/dt|\rangle\approx 1.363$, and $\langle|d\ell/dt|\rangle\approx 0.724$. Interestingly, although the trajectory approaches this ECS more closely than in the previous example, the dynamics are less synchronized with the ECS. Finally, we discuss a shadowing event in the interval $29944\tau,29947.2\tau]$, where $\langle d\rangle\approx 0.365$, $\langle|ds/dt|\rangle\approx 0.927$, and $\langle|d\ell/dt|\rangle\approx 0.267$. The trajectory approaches the ECS to a minimum distance of $d\approx 0.308$ at $t=29946.4$, which corresponds to one of the snapshots shown in the figure. Notice, however, that these shadowing confirmations are less pronounced than our findings in the more constrained attractor $\mathcal{X}$.
Figure 30: The dynamic shadowing criterion evaluated along the trajectory $Y(t)$ at discrete times (with $\delta t=0.4\tau$) within the interval $29000\tau,30000\tau]$. The plot marks instances where the dynamical similarity measure $\dfrac{ds}{dt}$ lies in the interval $[0.5,1.5]$, thereby satisfying the criterion. Although fewer shadowing events are detected than for $X(t)$, their presence points to similar underlying dynamics and motivates the search for additional ECSs to account for the remaining gaps.
Figure 31: Velocity field snapshots of three confirmed shadowing events for the trajectory $Y(t)$. The first row of every event shows three sequential snapshots (with $\delta t=0.4\tau$) of the turbulent trajectory. The second row of every event corresponds to sequential snapshots (at the same times) of the shadowed ECS.
The confirmed shadowing events suggest that the turbulence could follow paths that are broadly similar to those observed in the previous trajectory case. Specifically, we expect that the blue and red ECSs shape the nearly uniaxial flow, while the green ECSs mediate the transitions. However, as mentioned, the description with the ECS in hand remains incomplete. Consider, for example, the interval $29400\tau,29600\tau]$ (see Fig. 32 for a detailed view and Fig. 28 for the broader time series). Around $t=29412\tau$, the mean velocity $\langle u\rangle$ of the trajectory $Y(t)$ oscillates around $\langle u\rangle\approx 1.5$, showing an unclear shadowing of $\sigma_{y}$T1/13 with a distance $d\approx 0.585$. Following this, brief oscillations occur with successive approximations to the red, green, and blue ECSs. During this interval, there is a clear brief shadowing of the green ECS $\sigma_{x}$T1/3 at $t=29449.20$ with a minimum distance $d\approx 0.295$. The trajectory eventually reaches the longest and clearest oscillation in the blue region, where at $t=29537.6$, it satisfies the shadowing criterion for T1/13 with a distance $d\approx 0.787$. These three instants highlight the incomplete nature of our description: the large distances to the hypothetically shadowed ECSs when the trajectory is nearly unidirectional, along with the multiple ill-defined oscillations in between, illustrate the gaps in our current understanding.
Figure 32: Top row: Velocity field snapshots during the first reversal event for a turbulent trajectory $Y(t)$, evolving in the attractor $\mathcal{Y}(t)$, during the interval $29412\tau,29537.6\tau]$. Bottom row: Velocity field snapshots of the ECSs shadowed during this reversal, shown at the corresponding times. The resemblance between the turbulent flow and the shadowed ECS is noticeably weaker in snapshots where the flow is nearly unidirectional
Overall, the ECS based description of reversals is less clear here than in the case of $X(t)$. Nevertheless, the shadowing evidence indicates that ECSs and their invariant manifolds may still organize the dynamics within this attractor in a manner similar to what was observed for $X(t)$. A more comprehensive search for additional and potentially more intricate ECSs with longer time periods is therefore needed to resolve the gaps in our current understanding. Resolving these questions will be the subject of future work.
VI Discussion and Conclusions
Taking an infinite-dimensional dynamical systems perspective on active turbulence, we have studied a hydrodynamic model of active nematics in a minimal channel geometry. The dynamics in the smallest domain we consider is complex enough to sustain active turbulence, yet, as we have shown, simple enough that the phase space transport is organized around $\mathcal{O}(10)$ ECSs and the heteroclinic network connecting them.
Through an equivariant bifurcation analysis, we clarified the origin of several key solutions reported in earlier works and uncovered their role in shaping the global dynamics of the preturbulent regime. In this regime, three mechanisms for transient reversals are identified, involving either one of the two $y$-reflection-related copies of RPOs that are nearly homoclinic to the two uniaxial states, or the symmetric PO that is nearly heteroclinic to the two uniaxial states. Typical reversals involve a heteroclinic connection between one of the above mentioned ECSs (all unstable) and an attractor. Some reversals are mediated by a vortex-dominated PO. The homoclinic-like and heteroclinic-like ECSs possess the shift-reflect symmetry, and some of them additionally have discrete translation symmetries. All vortex-dominated POs have the $x$-$y$ reflection symmetry, with some again possessing discrete translation symmetries.
In the active turbulent regime in the original domain, we find that a subset of persistent flow reversals involve homoclinic-like RPOs that have slightly broken shift-reflect symmetry, but retain the discrete translation symmetry of their preturbulent counterparts. These reversals are mediated by vortex-dominated ECSs which are RPOs that have broken $x$-$y$ reflection symmetry yet still inherit the discrete translation symmetries of the preturbulent periodic orbits from which they originate. Building on this, we identify a smaller minimal flow unit corresponding to a fundamental domain [74], i.e., the smallest domain in which the lone turbulent attractor ($\mathcal{X}$) is organized by ECSs that no longer possess any residual spatial symmetries. In this domain, we establish robust shadowing of the dominating ECSs by typical turbulent trajectories via the use of symmetry-reduced distance and dynamical similarity measures.
Although our concrete computations are performed for a Beris–Edwards Q-tensor model, much of the structure we uncover relies only on symmetries that are shared across a broad class of continuum descriptions of active nematics. In particular, the local bifurcation analysis depends primarily on translational and reflection symmetries, and the nematic head–tail symmetry, and not on the precise forms of the equations [33]. The catalog of symmetry types of steady and time-periodic solutions and their bifurcations depends only on the total symmetry group $G$ and the representation on the critical eigenspaces, and therefore should be robust across all modeling choices that are expected to have these same symmetries [32]. We therefore expect the ECS families and reversal pathways identified here to be robust under variations in constitutive parameters and even under modest changes of the modeling framework. Repeating the presented analyses using different models will shift bifurcation points and stability boundaries, but the underlying symmetry-based organization and the existence of low-dimensional ECS skeletons mediating reversals should persist.
While shadowing of coherent structures is clearly observed and quantified in the smallest domain, conclusively establishing the same phenomenon in the original domain proved more challenging, owing to its richer dynamics, the larger dimension of the turbulent attractor, and our inability to compute a more complete set of ECSs. This difficulty is only expected to worsen in longer channels. Future work will focus on closing this gap and clarifying the role of invariant solutions in active nematic turbulence across a wider range of spatial and temporal scales. To do this, we will make use of recent advances in machine learning approaches [75, 76, 77] that have successfully employed autoencoders to compute ECSs in challenging inertial turbulent flow settings.
Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under award number DE-SC0022280. This work was completed utilizing the Holland Computing Center of the University of Nebraska, which receives support from the UNL Office of Research and Innovation, and the Nebraska Research Initiative.
Appendix
.1 Isotropy subgroups of unstable eigenspaces of UNI and LAN
Table 3: Isotropy groups of the unstable eigenspaces of UNI. All eigenvalues are complex. The $\mathrm{R}_{\mathrm{a}}$ column indicates the approximate activity values at which each two-dimensional unstable eigenspace emerges. From $\mathrm{R}_{\mathrm{a}}\approx 3.7$ to $4.5$, no additional local bifurcations from UNI are observed.
Complex pair Generators Group elements Group structure $\mathbf{R_{a}}$
1st $\{\sigma_{x}\,\tau_{x}(L/2)\}$ $\{e,\sigma_{x}\,\tau_{x}(L/2)\}$ $\mathbb{Z}_{2}$ 0.71
2nd $\{\sigma_{x}\,\tau_{x}(L/4)\}$ $\{e,\sigma_{x}\,\tau_{x}(L/4),\tau_{x}(L/2),\sigma_{x}\,\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 0.97
3rd $\{\sigma_{x},\tau_{x}(L/2)\}$ $\{e,\sigma_{x},\tau_{x}(L/2),\sigma_{x}\,\tau_{x}(L/2)\}$ $D_{2}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1.79
4th $\{\sigma_{x},\tau_{x}(L/3)\}$ $\{e,\tau_{x}(L/3),\tau_{x}(2L/3),\sigma_{x},\sigma_{x}\,\tau_{x}(L/3),\sigma_{x}\,\tau_{x}(2L/3)\}$ $\mathbb{Z}_{6}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{3}$ 2.03
5th $\{\sigma_{x}\,\tau_{x}(L/6)\}$ $\{e,\sigma_{x}\,\tau_{x}(L/6),\tau_{x}(L/3),\sigma_{x}\,\tau_{x}(L/2),$ $\tau_{x}(2L/3),\sigma_{x}\,\tau_{x}(5L/6)\}$ $\mathbb{Z}_{6}$ 2.5
6th $\{\sigma_{x}\,\tau_{x}(L/8)\}$ $\{e,\ \sigma_{x}\,\tau_{x}(L/8),\ \tau_{x}(L/4),\ \sigma_{x}\,\tau_{x}(3L/8),$ $\tau_{x}(L/2),\ \sigma_{x}\,\tau_{x}(5L/8),\ \tau_{x}(3L/4),\ \sigma_{x}\,\tau_{x}(7L/8)\}$ $\mathbb{Z}_{8}$ 3.01
7th $\{\sigma_{x},\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4),$ $\sigma_{x},\sigma_{x}\,\tau_{x}(L/4),\sigma_{x}\,\tau_{x}(L/2),\sigma_{x}\,\tau_{x}(3L/4)\}$ $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ 3.7
Table 4: Isotropy groups of the unstable eigenspaces of LAN. Each real eigenvalue has multiplicity two. The $\mathrm{R}_{\mathrm{a}}$ column indicates the approximate activity values at which each two-dimensional unstable eigenspace emerges.
Real pair Generators Group elements Group structure $\mathbf{R_{a}}$
1st $\{e\}$ $\{e\}$ Trivial group 0.85
2nd $\{\tau_{x}(L/2)\}$ $\{e,\,\tau_{x}(L/2)\}$ $\mathbb{Z}_{2}$ 1.25
3rd $\{\tau_{x}(L/2)\}$ $\{e,\tau_{x}(L/2)\}$ $\mathbb{Z}_{2}$ 2.11
4th $\{\tau_{x}(L/3)\}$ $\{e,\tau_{x}(L/3),\tau_{x}(2L/3)\}$ $\mathbb{Z}_{3}$ 2.51
5th $\{\tau_{x}(L/3)\}$ $\{e,\tau_{x}(L/3),\tau_{x}(2L/3)\}$ $\mathbb{Z}_{3}$ 3.10
6th $\{\tau_{x}(L/2),\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 3.30
7th $\{\tau_{x}(L/2),\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 4.21
8th $\{\tau_{x}(L/2),\tau_{x}(L/4)\}$ $\{e,\tau_{x}(L/4),\tau_{x}(L/2),\tau_{x}(3L/4)\}$ $\mathbb{Z}_{4}$ 4.36
.2 Symmetry classification: isotropy subgroups of the main ECS solutions
For ECSs in the preturbulent regime, the subscript is of the form $Tmx$, where the number $m$ labels the originating unstable eigenspace, while the letter x (=a, b, c, …) distinguishes different RPOs arising from the same eigenpair.
Each ECS in the turbulent regime is denoted $Tk/n$, where $k$ specifies the discrete translation symmetry $\tau_{x}(L_{A}/k)$, and $n$ indexes the different ECSs within that symmetry class.
Table 5: Main ECSs and their symmetries. All discrete translation symmetries refer to MFU A, with $L=L_{A}=\tilde{L}/3$.
Description Name Symmetry generator Comment
Homoclinic-like RPO
from 1st unstable pair of UNI HRPOT1a $\sigma_{x}\tau_{x}(L/2)$ Preturbulent regime.
Homoclinic-like RPO
from 2nd unstable pair of UNI HRPOT2a $\sigma_{x}\tau_{x}(L/4)$ Preturbulent regime.
Heteroclinic-like PO
from 1st unstable pair of UNI HTPOT1a $\sigma_{x}\tau_{x}(L/2)$ Preturbulent regime.
Heteroclinic-like PO
from 1st unstable pair of UNI HTPOT1b $\sigma_{x}\tau_{x}(L/2)$ Preturbulent regime.
One-Vortex PO
from LAN DPOT1a conjugated to $\sigma_{x}\sigma_{y}$ Preturbulent regime. Conjugation is by streamwise translation $\tau_{x}$.
RPOs possibly bifurcated
from DPO family DRPOT1a,b,c $\mathbb{I}$ Preturbulent regime.
Two-Vortex PO
from LAN POT2a conjugated to $\sigma_{x}\sigma_{y}\tau_{x}(L/2)$ Preturbulent regime. Conjugation is by streamwise translation $\tau_{x}$.
Homoclinic-like RPO
$T2/1,T2/15,T2/18,T2/7$
$\tau_{x}(L/2)$ Turbulent regime. Probable origin: HRPOT2a while breaking $\sigma_{x}\tau_{x}(L/4)$ symmetry
Vortex lattice RPO
$T2/9,T2/30,T2/4,T2/24,T2/13,T2/25$
$\tau_{x}(L/2)$ Turbulent regime. Probable origin: POT2a while breaking $\sigma_{x}\sigma_{y}$ symmetry
.3 Shadowing in Turbulent Regime: Additional information
Table 6: Top 10 ECS groups ranked by frequency of minimum distance occurrences of trajectory $X(t)$. These groups cover approximately 62% of all minimum distance instances.
ECS Group Freq. Rel. Freq. Cum. Freq. Cum. % Min Min Dist. Max Min Dist. Avg Min Dist. Region
T2/1 257 0.103 257 0.103 0.119 0.940 0.465 blue/red
T2/15 204 0.082 461 0.184 0.070 0.809 0.331 blue/red
T2/9 181 0.072 642 0.257 0.113 0.770 0.396 green
T2/7 161 0.064 803 0.321 0.256 1.132 0.561 blue/red
T2/18 134 0.054 937 0.375 0.139 1.056 0.513 blue/red
T2/30 134 0.054 1071 0.428 0.203 0.607 0.341 green
T2/4 133 0.053 1204 0.481 0.154 0.761 0.329 green
T2/24 118 0.047 1322 0.529 0.244 0.648 0.394 green
T2/13 116 0.046 1438 0.575 0.207 0.835 0.396 green
T2/25 110 0.044 1548 0.619 0.153 0.807 0.412 green
Table 7: Top 15 ECS groups ranked by frequency of minimum distance occurrences of trajectory Y(t). These groups cover approximately 62% of all ‘closest approach’ instances.
Trajectory ID Freq. Rel. Freq. Cum. Freq. Cum. % Min Min Dist. Max Min Dist. Avg Min Dist. Region
T1/6 167 0.067 167 0.067 0.308 1.046 0.593 blue/red
T1/3 158 0.063 325 0.130 0.243 1.011 0.611 green
T1/11 158 0.063 483 0.193 0.306 0.996 0.678 blue/red
T1/18 144 0.058 627 0.251 0.399 1.027 0.662 blue/red
T1/23 113 0.045 740 0.296 0.572 1.039 0.790 green
T1/24 108 0.043 848 0.339 0.533 1.010 0.742 green
T1/17 101 0.040 949 0.379 0.465 0.965 0.694 blue/red
T1/2 99 0.040 1048 0.419 0.393 1.009 0.660 green
T1/37 85 0.034 1133 0.453 0.299 0.955 0.652 blue/red
T1/21 78 0.031 1211 0.484 0.556 1.011 0.782 green
T1/25 77 0.031 1288 0.515 0.489 1.036 0.724 green
T1/13 74 0.030 1362 0.545 0.539 1.104 0.772 green
T1/33 71 0.028 1433 0.573 0.424 0.879 0.624 blue/red
T1/29 61 0.024 1494 0.597 0.596 0.927 0.742 green
T1/27 59 0.024 1553 0.621 0.320 0.870 0.610 blue/red
Table 8: PCA results comparing the two trajectories embedded in chaotic sets $\mathcal{X}$ and $\mathcal{Y}$, respectively
PCA Metric $X(t)$ $Y(t)$
Number of time steps analyzed 2501 2501
Number of components explaining 90% variance 4 8
Number of components explaining 95% variance 6 12
Total number of components (time points) 2501 2501
Variance explained by PC1 (%) 60.37 28.41
Variance explained by PC2 (%) 24.40 18.11
Variance explained by PC3 (%) 4.28 16.01
Variance explained by PC4 (%) 4.13 15.32
Variance explained by PC5 (%) 1.61 3.62
Variance explained by PC6 (%) 1.25 3.05